### v1

# Cluster-cyclic condition of skew-symmetrizable matrices of rank 3 via the Markov constant

Abstract.
In this paper, we consider mutations of skew-symmetrizable matrices of rank 3. Every skew-symmetrizable matrix corresponds to a weighted quiver, and we study the conditions when this quiver is always cyclic after applying mutations. In this study, the Markov constant has an essential meaning. It has already appeared in some previous works for skew-symmetric matrices.
1. Introduction
1.1. Background
The cluster algebra theorey was introduced by Fomin and Zelevinsky in [FZ02]. The main objects are mutations, which are transformations of matrices and rational fractions called cluster variables. In this paper, we consider mutations of skew-symmetrizable matrices of rank 3. The matrix $B\in\mathrm{M}_{3}(\mathbb{R})$ is called skew-symmetrizable if there exists a diagonal matrix $D$ with positive diagonal entries such that $DB$ is skew-symmetric. Then, it was shown in [FZ03] that $B$ may be expressed as
(1) $B=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\end{matrix}\right)$
where $x,y,z,x^{\prime},y^{\prime},z^{\prime}\in\mathbb{R}$ satisfy $xyz=x^{\prime}y^{\prime}z^{\prime}$. In the usual cluster algebra theory, we often suppose that all entries of $B$ are integral, since entries of $B$ appear in the exponent of rational fractions. However, since we do not focus on cluster variables, we consider real skew-symmetrizable matrices. When we focus on only matrix mutations $\mu_{k}(B)$, we can naturally extend the definition, and their properties are helpful for the study of cluster algebras. This generalization for real entries have already appeared in [BBH11, FT19, FT23] for the skew-symmetric case.
The rules of mutations depend on the signs of entries in the matrix. If all $x,x^{\prime},y,y^{\prime},z$, and $z^{\prime}$ are either positive or negative, then $B$ is called cyclic. Otherwise, $B$ is called acyclic. This word comes from the definitions in the corresponding quivers $Q(B)$. (See Figure 2 and Figure 2.) If every matrix obtained by applying mutations to $B$ is cyclic, then $B$ is called cluster-cyclic, and otherwise, we call $B$ cluster-acyclic. Our purpus is to identify whether $B$ is cluster-cyclic or not.
In the skew-symmetric case, this problem was solved in [BBH11, FT19]. In [BBH11], they showed it for the integer skew-symmetric case, and in [FT19], they generalized it for real skew-symmetric case by using the geometric aspect of mutations. They showed that the Markov constants, which was introduced in [BBH11] for the cluster algebra theory, play important roles in this problem. For the skew-symmetrizable case, the Markov constants $C(B)=xx^{\prime}+yy^{\prime}+zz^{\prime}-|xyz|$, which are a generalization of the one for skew-symmetric matrices, also play important roles in this problem.
Other motivation of this study is in the classification of the global structures of $G$-fans [Nak24].
1.2. Method
To study the cluster-cyclicity, we combine the method of [FZ03] and [BBH11]. In [FZ03], there is a map from skew-symmetrizable matrices to skew-symmetric matrices, which we write $\mathrm{Sk}(B)$ in this paper, such that this operation $\mathrm{Sk}$ and mutations $\mu_{k}$ are commutative. (See Proposition 2.6.) By using this correspondence, some methods for skew-symmetric matrices in [BBH11] work well for skew-symmetrizable matrices. By translating the statements of skew-symmetrizable matrices into the ones of skew-symmetric matrices, some results have already been shown in [FT19]. In particular, Theorem 1.1 below is immediately shown. However, since we use the method of the proof in [BBH11] in Section 9 and Section 10, we give another proof which is similar to the method of [BBH11].
However, by considering skew-symetrizable matrices, we need some reconstructions. One of the main reasons is that the surjectivity of Markov constants. (See Theorem 10.12, Example 10.18.) In [BBH11], it was shown that there are no integer skew-symmetric and cluster-cyclic matrices with $40<C(B)<$ 4. Thanks to this property, some arguments became clear. However, by considering skew-symmetrizable matrices, there are both cluster-cyclic and cluster-acyclic matrices $B$ with $40<C(B)<$ 4.
1.3. Main results
First, we mention the identification of cluster-cyclicity. It is a natural generalization of [BBH11] for the skew-symmetrizable and real case.
Theorem 1.1 (cf. [BBH11, Thm. 1.2 (1),(2)]).
Let $B$ be a skew-symmetrizable and cyclic matrix of the form (1). Then, the following two conditions are equivalent.
  • 1.
    $B$ is cluster-cyclic.
  • 2.
    It holds that $C(B)\leq 4$ and $xx^{\prime},yy^{\prime},zz^{\prime}\geq 4$.
For some thechnical reasons, it is better to consider transformations $\gamma_{k}=-\mu_{k}$ than the usual mutations $\mu_{k}$ for any $k=1,2,3$. Let $\Gamma$ be a group generated by $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$. When $B$ of the form (1) satisfies $x,y,z,x^{\prime},y^{\prime},z^{\prime}>0$, $B$ is called positive-cyclic. By considering $\gamma_{k}$, if $B$ is positive and cluster-cyclic, then every element in this orbit $\Gamma(B)$ is also so. (See Proposition 4.1.)
Next result is about a fundamental domain for this action $\Gamma$ of integer skew-symmetrizable matrices. The following partial order $\leq$ on skew-symmetrizable and positive-cyclic matrices is useful in this study.
(2) $\ \left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\end{matrix}\right)\leq\left(\begin{matrix}0&-w^{\prime}&v\\ w&0&-u^{\prime}\\ -v^{\prime}&u&0\end{matrix}\right)$
$\Longleftrightarrow$ $\ x\leq u,\ y\leq v,\ z\leq w,\ x^{\prime}\leq u^{\prime},\ y^{\prime}\leq v^{\prime},\ z^{\prime}\leq w^{\prime}.$
By using this order, we can give a fundamental domain for $\Gamma$, which is also a natural generalization of [BBH11, Thm. 5.1].
Theorem 1.2 (Theorem 10.7).
A set $\tilde{F}$ defined by
(3) $\left\{B=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\\ \end{matrix}\right)\mid\begin{aligned} &\textup{$B$ is skew-symmetrizable},\\ &xyz\geq 2xx^{\prime},2yy^{\prime},2zz^{\prime},\\ &x,y,z,x^{\prime},y^{\prime},z^{\prime}\in\mathbb{Z}_{>0}\end{aligned}\right\}$
is a fundamental domain for $\Gamma$ on integer skew-symmetrizable, positive and cluster-cyclic matrices. Moreover, for any integer skew-symmetrizable, positive, and cluster-cyclic matrix $B$, the element of $\Gamma(B)\cap\tilde{F}$ is the smallest element in $\Gamma(B)$.
Nemely, for each integer skew-symmetrizable, positive, and cluster-cyclic matrix $B$, there is the minimum element of $\Gamma(B)$ and it is an element in $\tilde{F}$. Unfortunately, this result cannnot generalize for the real case. (See Theorem 9.1.)
These two results are natural generalizations of the skew-symmetric case. On the other hand, the following result is the most important difference between the skew-symmetrizable case and the skew-symmetric case.
Theorem 1.3 (Theorem 10.12).
The Markov constant $C(B)$ can take any integer less than or equal to $4$ by choosing an integer skew-symmetrizable and cluster-cyclic matrix $B$.
Because of this property (in particular, the fact that there is a cluster-cyclic matrix $B$ with $40<C(B)<$ 4), we need to reconstruct a proof in [BBH11] even though we focus on the integer skew-symmetrizable matrices (see Example 10.18).
1.4. Structure of the paper
In section 2, we recall the properties of mutations introduced by [FZ02, FZ03].
In section 3, we define the Markov constant. This is a generalization for the skew-symmetrizable case of the one which is introduced in [BBH11].
In section 4, we introduce more simple notations, which is similar to [BBH11, FT19]. In this notations, Theorem 1.1 is decomposed to Lemma 4.7, Lemma 4.8, and Lemma 4.9.
In section 5, we show Lemma 4.7.
In section 6, we calculate some examples and prove Lemma 4.8. Most contents and ideas were in [FT19].
In section 7 and 8, we show Lemma 4.9. Most ideas are similar to the one in [BBH11]. However some modifications are necessarily to generalize it for the real and skew-symmetrizable case.
In section 9 and 10, we derive some new results. In section 9, we consider the case which does not occur when we consider integer skew-symmetric matrces. The related topic is in [FT19]. In section 10, we deal with the integer skew-symmetrizable case. In particular, we focus on the difference between the result of the skew-symmetric case in [BBH11] and the one of the skew-symmetrizable case.
Acknowledgement.
This problem was suggested by Tomoki Nakanishi, and I thank him for various important advices. I also thank Zhichao Chen for careful reading and advices.
This work was financially supported by JST SPRING, Grant Number JPMJSP2125. The author would like to take this opportunity to thank the “THERS Make New Standards Program for the Next Generation Researchers.”
2. Definitions
2.1. Skew-symmetrizable matrix and mutation
In this paper, we will focus on skew-symmetrizable matrices of rank $3$ with real entries. Let $\mathrm{M}_{3}(\mathbb{R})$ be the set of all $3\times 3$ real matrices. A real matrix $B=(b_{ij})\in\mathrm{M}_{3}(\mathbb{R})$ is called skew-symmetrizable if there exists a diagonal matrix $D=\mathrm{diag}(d_{1},d_{2},d_{3})$ with positive real numbers $d_{1},d_{2},d_{3}>0$ such that $DB$ is skew-symmetric. This $D$ is said to be a skew-symmetrizer of $B$.
Skew-symmetrizable matrices have the following properties.
Proposition 2.1 ([FZ03, Lem. 7.4]).
Let $B=(b_{ij})\in\mathrm{M}_{3}(\mathbb{R})$ be a skew-symmetrizable matrix. Then, we have the following statements.
(a) $B$ is sign-skew-symmetric, namely, $b_{ij}$ and $-b_{ji}$ have the same sign for any $i,j$. (In particular, every diagonal entry is 00.)
(b) We have
(4) $b_{12}b_{23}b_{31}=-b_{21}b_{32}b_{13}.$
Conventionally, the sign $\varepsilon_{x}$ of a real number $x\in\mathbb{R}$ is defined as follows: $\varepsilon_{x}=0$ if $x=0$ and, otherwise, $\varepsilon_{x}=\pm 1$ with $\varepsilon_{x}x>0$. Next, we difine mutations of skew-symmetrizable matrices.
Definition 2.2.
Let $B=(b_{i,j})\in\mathrm{M}_{3}(\mathbb{R})$ be a skew-symmetrizable matrix. We define mutations $\mu_{k}(B)=(b^{\prime}_{ij})\in\mathrm{M}_{3}(\mathbb{R})$ ($k=1,2,3$) as follows:
(5) $b^{\prime}_{ij}=\begin{cases}-b_{ij}&i=k\ \textup{or}\ j=k,\\ b_{ij}+b_{ik}[b_{kj}]_{+}+[-b_{ik}]_{+}b_{kj}&i,j\neq k,\end{cases}$
where $[a]_{+}$ is defined as $[a]_{+}=\max(a,0)$.
It is known that $\mu_{k}\mu_{k}(B)=B$ holds for any $B\in\mathrm{M}_{3}(\mathbb{R})$ and $k=1,2,3$ [FZ02]. We can check that every skew-symmetrizer of $B$ is also a skew-symmetrizer of $\mu_{k}(B)$.
For any skew-symmetrizable matrix $B$, we may define the corresponding quiver $Q(B)$ wighted by $\mathbb{R}_{\geq 0}^{2}$ as follows:
  • The vertices of $Q(B)$ are indexed by $1$, $2$, and $3$.
  • For each vertex pair $i\neq j$, if $b_{ij}>0$ or, equivalently, $b_{ji}<0$, there is an edge $i\overset{(b_{ij},-b_{ji})}{\longrightarrow}j$. If $b_{ij}<0$ or $b_{ji}>0$, there is an edge $j\overset{(b_{ji},-b_{ij})}{\longrightarrow}i$. If $b_{ij}=b_{ji}=0$, there are no edges between $i$ and $j$.
A skew-symmetrizable matrix $B$ is called cyclic if the corresponding quiver $Q(B)$ is cyclic. Otherwise, $B$ is said to be acyclic. By definition, $B$ is cyclic if and only if $B$ may be expressed as
(6) $\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\end{matrix}\right)$
for some $x,y,z,x^{\prime},y^{\prime},z^{\prime}\in\mathbb{R}$ such that all of these numbers are either positive or negative. When all of these numbers are positive (resp. negative), this matrix is said to be positive-cyclic (resp. negative-cyclic). The corresponding quivers are described as in Figure 2 and Figure 2.
111122223333$(y,y^{\prime})$$(x,x^{\prime})$$(z,z^{\prime})$
Figure 1. Positive
111122223333$(-y^{\prime},-y)$$(-x^{\prime},-x)$$(-z^{\prime},-z)$
Figure 2. Negative
Next, we define the cluster-cyclicity. This is the main contents in this paper.
Definition 2.3.
Let $B\in\mathrm{M}_{3}(\mathbb{R})$ be a skew-symmetrizable matrix. If every matrix which is obtained by mutating $B$ finitely many times is cyclic, then $B$ is said to be cluster-cyclic; otherwise, $B$ is said to be cluster-acyclic.
For any skew-symmetrizable matrix $B$, if $B$ is positive-cyclic (resp. negative-cyclic), then $\mu_{k}(B)$ is not positive-cyclic because every entry in the $k$-th row or column changes their signs. In particular, if $B$ is cluster-cyclic, positive-cyclic and negative-cyclic matrices are generated alternately by applying mutations.
2.2. Double-sided-skew-symmetrized form
Our subsequent purpus is to identify whether a given skew-symmetrizable matrix $B\in\mathrm{M}_{3}(\mathbb{R})$ is cluster-cyclic or not. For this, it is enough to consider the certain skew-symmetric matrix, which we name a double-sided-skew-symmetrized form. It was introduced by [FZ03, Lem. 8.3] when they consider the mutations of diagrams corresponding to skew-symmetrizable matrices.
For any real number $x$, we write the sign of $x$ by $\varepsilon_{x}\in\{0,\pm 1\}$.
Definition 2.4.
Let
(7) $B=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\end{matrix}\right)\in\mathrm{M}_{3}(\mathbb{R})$
be a (not necessarily cyclic) skew-symmetrizable matrix. (Note that $x$, $y$, and $z$ have the same sign as $x^{\prime}$, $y^{\prime}$, and $z^{\prime}$, respectively.) Set $p=\varepsilon_{x}\sqrt{xx^{\prime}}$$q=\varepsilon_{y}\sqrt{yy^{\prime}}$, and $r=\varepsilon_{z}\sqrt{zz^{\prime}}$. Then, define the skew-symmetric matrix $\mathrm{Sk}(B)\in\mathrm{M}_{3}(\mathbb{R})$ as
(8) $\mathrm{Sk}(B)=\left(\begin{matrix}0&-r&q\\ r&0&-p\\ -q&p&0\end{matrix}\right).$
We call it the double-sided-skew-symmetrized form of $B$.
The name of double-sided-skew-symmetrized comes from the following property.
Lemma 2.5 ([FZ03, Lem. 8.3]).
Let $B\in\mathrm{M}_{3}(\mathbb{R})$ be a skew-symmetrizable matrix. Then, for arbitrary skew-symmetrizer $D=\mathrm{diag}(d_{1},d_{2},d_{3})$ of $B$, by choosing its square root $D^{1/2}=\mathrm{diag}(\sqrt{d_{1}},\sqrt{d_{2}},\sqrt{d_{3}})$, it holds that
(9) $D^{\frac{1}{2}}BD^{-\frac{1}{2}}=\mathrm{Sk}(B),$
where $D^{-1/2}=(D^{1/2})^{-1}=\mathrm{diag}\left(\sqrt{d_{1}}^{-1},\sqrt{d_{2}}^{-1},\sqrt{d_{3}}^{-1}\right)$.
In the notations of Definition 2.4, we often use the following equality.
(10) $xyz=x^{\prime}y^{\prime}z^{\prime}=pqr.$
The first equality follows from Proposition 2.1 (b). The second equality may be verified as follows:
(11) $xyz=\varepsilon_{xyz}\sqrt{|xyz|}\sqrt{|x^{\prime}y^{\prime}z^{\prime}|}=\varepsilon_{x}\varepsilon_{y}\varepsilon_{z}\sqrt{|xx^{\prime}|}\sqrt{|yy^{\prime}|}\sqrt{|zz^{\prime}|}=pqr.$
Note that $|xyz|=|x^{\prime}y^{\prime}z^{\prime}|=\sqrt{|xyz|}\sqrt{|x^{\prime}y^{\prime}z^{\prime}|}$.
This double-sided-skew-symmetrized form is compatible for the cluster-cyclicity.
Proposition 2.6.
Let $B\in\mathrm{M}_{3}(\mathbb{R})$ be a skew-symmetrizable matrix. Then, the following statements hold:
(a) Every entry of $\mathrm{Sk}(B)$ has the same sign as the corresponding entry of $B$. In patricular, $B$ is cyclic if and only if $\mathrm{Sk}(B)$ is cyclic.
(b)[FZ03, Lem. 8.4] For any $k=1,2,3$, we have
(12) $\mu_{k}(\mathrm{Sk}(B))=\mathrm{Sk}(\mu_{k}(B)).$
In particular, $B$ is cluster-cyclic if and only if $\mathrm{Sk}(B)$ is cluster-cyclic.
The statement (a) is obvious by the definition, and (b) have been shown in [FZ03]. By combining (a) and (b), we can show that the last staement by the induction on the number of mutations.
3. Markov constant
In this section, we define the Markov constant of a skew-symmetrizable matrix. In [BBH11], it has an important role in the problem for the integer skew-symmetric case. It is also important for the skew-symmetrizable case.
In this section, we fix a skew-symmetrizable matrix and its double-sided-skew-symmetrized form as
(13) $B=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\end{matrix}\right),\ \mathrm{Sk}(B)=\left(\begin{matrix}0&-r&q\\ r&0&-p\\ -q&p&0\end{matrix}\right).$
as in Definition 2.4. By Proposition 2.1 and (10), these entries satisfy the following conditions:
  • In each triplet $(x,x^{\prime},p)$, $(y,y^{\prime},q)$, and $(z,z^{\prime},r)$, the three entries have the same sign.
  • It holds that $xyz=x^{\prime}y^{\prime}z^{\prime}=pqr$.
The following definition is suggested by Tomoki Nakanishi.
Definition 3.1.
For any skew-symmetrizable matrix $B\in\mathrm{M}_{3}(\mathbb{R})$ of the form (13), we define the Markov constant $C(B)$ of $B$ as follows:
(14) $C(B)=xx^{\prime}+yy^{\prime}+zz^{\prime}-|xyz|.$
Similarly, we define $C^{+}(B)$ and $C^{-}(B)$ as
(15) $C^{\pm}(B)=xx^{\prime}+yy^{\prime}+zz^{\prime}-(\pm xyz).$
These three numbers have the following relations.
Proposition 3.2.
Let $B$ be a skew-symmetrizable matrix.
(a) We have $C(B)=C(-B)$ and $C^{+}(B)=C^{-}(-B)$.
(b) If $B$ is positive-cyclic (resp. negative-cyclic), then we have $C(B)=C^{+}(B)$ (resp. $C(B)=C^{-}(B)$).
We may obtain these properties by considering a sign of $xyz$.
For any skew-symmetrizable matrix $B$, since $xyz=x^{\prime}y^{\prime}z^{\prime}=pqr$ as shown in (10), the Markov constant $C(B)$ may also be expressed as follows:
(16) $C(B)$ $=xx^{\prime}+yy^{\prime}+zz^{\prime}-|xyz|$
$=xx^{\prime}+yy^{\prime}+zz^{\prime}-|x^{\prime}y^{\prime}z^{\prime}|$
$=p^{2}+q^{2}+r^{2}-|pqr|=C(\mathrm{Sk}(B)).$
The following proposition is fundamental for later.
Proposition 3.3.
Let $B$ be a skew-symmetrizable matrix. Then, the following two statements hold.
(a) We have $C^{\pm}(B)=C^{\pm}(\mathrm{Sk}(B))$ and $C(B)=C(\mathrm{Sk}(B))$.
(b) If $B$ is positive-cyclic (resp. negative-cyclic), we have $C^{-}(\mu_{k}(B))=C^{+}(B)$ (resp. $C^{+}(\mu_{k}(B))=C^{-}(B)$) for any $k=1,2,3$. In particular, we have $C^{+}(-\mu_{k}(B))=C^{+}(B)$ if $B$ is positive-cyclic.
Proof.
Let $B$ be expressed as in (13). The statement (a) has already been shown in (16). We show (b) in the case of $k=1$. Let $\varepsilon$ be $+$ and -- if $B$ is positive-cyclic and negative-cyclic, respectively. Since $C^{\varepsilon}(B)=C^{\varepsilon}(\mathrm{Sk}(B))$ and $C^{-\varepsilon}(\mu_{1}(B))=C^{-\varepsilon}(\mathrm{Sk}(\mu_{1}(B)))=C^{-\varepsilon}(\mu_{1}(\mathrm{Sk}(B)))$ by (a) and Proposition 2.6 (b), it suffices to show that $C^{\varepsilon}(\mathrm{Sk}(B))=C^{-\varepsilon}(\mu_{1}(\mathrm{Sk}(B)))$. We have
(17) $\mu_{1}(\mathrm{Sk}(B))=\left(\begin{matrix}0&r&-q\\ -r&0&-p+\varepsilon qr\\ q&p-\varepsilon qr&0\end{matrix}\right).$
Thus, we obtain
(18) $C^{-\varepsilon}(\mu_{1}(\mathrm{Sk}(B)))$ $=(p-\varepsilon qr)^{2}+q^{2}+r^{2}-(-\varepsilon)(p-\varepsilon qr)(-q)(-r)$
$=p^{2}+q^{2}+r^{2}-\varepsilon pqr=C^{\varepsilon}(\mathrm{Sk}(B)).$
4. Reduction of Theorem 1.1 to the skew-symmetric case
In this section, we give more simple notations as in [BBH11], and we translate Theorem 1.1 into these notations.
Let $B\in\mathrm{M}_{3}(\mathbb{R})$ be a skew-symmetrizable matrix. Let $t=(t_{1},t_{2},\dots,t_{l})$ be a finite sequence of indices $1,2,3$ such that $t_{i}\neq t_{i+1}$ for any $i=1,2,\dots,l-1$. The length $|t|$ is defined by the number of indices; in this case, the length of $t$ is $l$. We write $\mu_{t}(B)=\mu_{t_{l}}\cdots\mu_{t_{2}}\mu_{t_{1}}(B)$. A collection $\Sigma(B)=\{\mu_{t}(B)\}_{t}$ indexed by all of such $t$ is called the matrix-pattern of $B$. In convention, we allow $t=\emptyset$ and let $\mu_{\emptyset}=\mathrm{id}$. We often fix one element of $\Sigma(B)$ and we call it an initial matrix of $\Sigma(B)$. Now, we may see $B$ as an initial marix of $\Sigma(B)$. Then the distance (from $B$) to $\mu_{t}(B)\in\Sigma(B)$ is defined by the length of $t$.
Let
(19) $B=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\end{matrix}\right)$
be a skew-symmetrizable and cyclic matrix. Recall that $B$ is said to be positive-cyclic (resp. negative-cyclic) if all of $x,y,z,x^{\prime},y^{\prime},z^{\prime}$ are positive (resp. negative). Later, we often consider positive-cyclic matrices. If we consider a negative-cyclic matrix $B$, apply the results for $-B$. Our interest is whether a given positive-cyclic matrix $B\in\mathrm{M}_{3}(\mathbb{Z})$ is cluster-cyclic or not. In this case, it is enough to consider the following maps $\gamma_{k}$ instead of $\mu_{k}$.
(20) $\gamma_{1}(B)$ $=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-(yz-x^{\prime})\\ -y^{\prime}&y^{\prime}z^{\prime}-x&0\end{matrix}\right),$
$\gamma_{2}(B)$ $=\left(\begin{matrix}0&-z^{\prime}&z^{\prime}x^{\prime}-y\\ z&0&-x^{\prime}\\ -(zx-y^{\prime})&x&0\end{matrix}\right),$
$\gamma_{3}(B)$ $=\left(\begin{matrix}0&-(xy-z^{\prime})&y\\ x^{\prime}y^{\prime}-z&0&-x^{\prime}\\ -y^{\prime}&x&0\end{matrix}\right),$
where the domain of these functions consists of all skew-symmetrizable matrices (not necessarily cyclic).
We may verify that $\gamma_{k}\gamma_{k}(B)=B$. Define $\Gamma(B)=\{\gamma_{t}(B)\}_{t}$ in the same way as $\Sigma(B)$, but we replace $\mu_{k}$ with $\gamma_{k}$. In this case, we have the following proposition.
Proposition 4.1.
Let $B$ be a skew-symmetrizable and positive-cyclic matrix.
(a) We have $\gamma_{k}(B)=-\mu_{k}(B)$ for any $k=1,2,3$.
(b) The following two conditions are equivalent.
  • (1)
    $B$ is cluster-cyclic.
  • (2)
    Every matrix of $\Gamma(B)$ is positive-cyclic.
Proof.
We may show (a) by the direct calculation. By considering (a), we may show (b) as follows: Note that $\mu_{k}(-B)=-\mu_{k}(B)$. For a sequence of indices $t=(t_{1},\dots,t_{l})$ and $t^{\prime}=(t_{1},\dots,t_{l},t_{l+1})$, if $\gamma_{t}(B)=(-1)^{|t|}\mu_{t}(B)$ is cluster-cyclic, we have $\gamma_{t^{\prime}}(B)=(-1)^{|t^{\prime}|}\mu_{t^{\prime}}(B)$. This relation is true when $\mu_{t}(B)$ is cyclic. (we can disregard the condition of whether $\mu_{t^{\prime}}(B)$ is cyclic or not.) Thus, if there is an acyclic-matrix $\mu_{t^{\prime}}(B)$ and $|t^{\prime}|$ is the smallest among such $t^{\prime}$, then it holds that $\gamma_{t^{\prime}}(B)=(-1)^{|t^{\prime}|}\mu_{t^{\prime}}(B)$ and it is acyclic. ∎
Later, we write a skew-symmetrizable matrix by a following tuplet:
(21) $M=\left(\begin{matrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{matrix}\right).$
By Proposition 2.1, these numbers satisfy the following properties.
(22) The equality $xyz=x^{\prime}y^{\prime}z^{\prime}$ holds
and the sign of two entries on the same row is the same.
The maps $\gamma_{k}$ may be expressed as
(23) $\gamma_{1}(M)$ $=\left(\begin{matrix}y^{\prime}z^{\prime}-x&y&z\\ yz-x^{\prime}&y^{\prime}&z^{\prime}\end{matrix}\right),$
$\gamma_{2}(M)$ $=\left(\begin{matrix}x&z^{\prime}x^{\prime}-y&z\\ x^{\prime}&zx-y^{\prime}&z^{\prime}\end{matrix}\right),$
$\gamma_{3}(M)$ $=\left(\begin{matrix}x&y&x^{\prime}y^{\prime}-z\\ x^{\prime}&y^{\prime}&xy-z^{\prime}\end{matrix}\right).$
Moreover, by Proposition 2.6 (c), when we consider whether a skew-symmetrizable matrix $B\in\mathrm{M}_{3}(\mathbb{R})$ is cluster-cyclic or not, it suffices to consider whether
(24) $\mathrm{Sk}(B)=\left(\begin{matrix}0&-r&q\\ r&0&-p\\ -q&p&0\end{matrix}\right)$
is cluster-cyclic or not. We may see this matrix as a triplet of real numbers $(p,q,r)\in\mathbb{R}^{3}$ and the maps $\gamma_{k}$ become
(25) $\gamma_{1}(p,q,r)$ $=(qr-p,q,r),$
$\gamma_{2}(p,q,r)$ $=(p,rp-q,r),$
$\gamma_{3}(p,q,r)$ $=(p,q,pq-r).$
We summerize the definitions.
Definition 4.2.
Define sets
[…]
+=$\mathcal{M}^{+}=\left\{\left(\begin{matrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{matrix}\right)\in\mathcal{M}\mid x,y,z,x^{\prime},y^{\prime},z^{\prime}>0\right\},$
$\mathcal{S}=\{(p,q,r)\mid p,q,r\in\mathbb{R}\}=\mathbb{R}^{3},$
$\mathcal{S}^{+}=\{(p,q,r)\in\mathcal{S}\mid p,q,r>0\}.$
Every element of $\mathcal{M}^{+}$ and $\mathcal{S}^{+}$ is called positive. We define maps $\gamma_{k}:\mathcal{M}\to\mathcal{M}$ and $\gamma_{k}:\mathcal{S}\to\mathcal{S}$ as (23) and (25), respectively.
The maps $\gamma_{k}$ satisfy $\gamma_{k}^{2}=\mathrm{id}$, in particular, $\gamma_{k}$ is a bijection on $\mathcal{S}$.
Definition 4.3.
For any $M\in\mathcal{M}$, let $\Gamma(M)$ be the collection of $\gamma_{t}(M)$ indexed by all of finite sequences $t=(t_{1},\dots,t_{k})$ of indices $1,2,3$ with $t_{i}\neq t_{i+1}$, where $\gamma_{t}=\gamma_{t_{k}}\cdots\gamma_{t_{1}}$. For any positive element $M\in\mathcal{M}^{+}$, if every element in $\Gamma(M)$ is positive, $M$ is said to be cluster-positive. Similarly, we define the cluster-positivity of $S\in\mathcal{S}^{+}$.
As we said, every positive element $M\in\mathcal{M}^{+}$ and $S\in\mathcal{S}^{+}$ corresponds to a positive-cyclic matrices. Here, we define a map $\mathrm{Sk}:\mathcal{M}\to\mathcal{S}$ as
(27) $\mathrm{Sk}\left(\begin{matrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{matrix}\right)=(\varepsilon_{x}\sqrt{xx^{\prime}},\varepsilon_{y}\sqrt{yy^{\prime}},\varepsilon_{z}\sqrt{zz^{\prime}}),$
where $\varepsilon_{a}\in\{0,\pm 1\}$ is the sign of $a\in\mathbb{R}$. Obviously, if $M\in\mathcal{M}$ corresponds to a positive-cyclic matrix $B\in\mathrm{M}_{3}(\mathbb{R})$, then $\mathrm{Sk}(M)$ corresponds to its double-sided skew-symmetrized form $\mathrm{Sk}(B)$. Moreover, $M$ is cluster-positive if and only if $S=\mathrm{Sk}(M)$ is cluster-positive.
Definition 4.4.
For any $M=\left(\begin{smallmatrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{smallmatrix}\right)\in\mathcal{M}$ and $S=(p,q,r)\in\mathcal{S}$, we define the Markov constants $C(M)=xx^{\prime}+yy^{\prime}+zz^{\prime}-xyz$ and $C(S)=p^{2}+q^{2}+r^{2}-pqr$.
By definition, we have $C(\mathrm{Sk}(M))=C(M)$. Note that, for any positive-cyclic and skew-symmetrizable matrix $B$, it corresponds to an element $M\in\mathcal{M}^{+}$. Then, it holds that
(28) $C(M)=C^{+}(B)=C(B).$
Moreover, we have
(29) $C(\gamma_{k}(M))=C(M).$
Proof is similar to the one of Proposition 3.3 (b), in particular, (18). Thanks to this result, the Markov constant does not change under the transformations of $\gamma_{k}$.
We summarize the results in this section.
Proposition 4.5.
Let
(30) $B=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\end{matrix}\right)\in\mathrm{M}_{3}(\mathbb{R})$
be skew-symmetrizable and positive-cyclic. Set $p=\sqrt{xx^{\prime}}$, $q=\sqrt{yy^{\prime}}$, $r=\sqrt{zz^{\prime}}$, and $S=(p,q,r)\in\mathcal{S}^{+}$.
(Proposition 4.1) $B$ is cluster-cyclic if and only if $S$ is cluster-cyclic.
(28) We have $C(B)=C(S)$.
(29) On $\mathcal{S}$, the Markov constant does not change under the transformations of $\gamma_{k}$.
Based on these results, we may translate Theorem 1.1 into these notations. This is the generalization of [BBH11, Thm. 1.2 (1),(2)], and it was shown in [FT19].
Proposition 4.6 ([FT19, Lem. 3.3, Thm. 4.5, Cor. 4.8]).
Let $S=(p,q,r)\in\mathcal{S}^{+}$. Then, $S$ is cluster-cyclic if and only if both $p,q,r\geq 2$ and $C(S)\leq 4$ hold.
Now, we show Theorem 1.1 under the assumption of Proposition 4.6.
Proof of Theorem 1.1.
Let $B$ be a skew-symmetrizable and cyclic matrix. If $B$ is positive-cyclic, the claim holds by using the correspondence of $\mathrm{Sk}(B)$ and $S$. When $B$ is negative-cyclic, then $-B$ is positive-cyclic. Thus, by applying the result of Proposition 4.6 for $-B$, we can show Theorem 1.1. ∎
Proposition 4.6 was shown in [FT19] by using geometrical aspects of mutations. (Thus, we complete the proof of Theorem 1.1.) However, we give an alternatively proof of this proposition, which is more closely related to [BBH11].
We decompose Proposition 4.6 into the following three statemenets.
Lemma 4.7.
Let $S=(p,q,r)\in\mathcal{S}^{+}$ satisfy $C(S)\leq 4$ and $p,q,r\geq 2$. Then, $S$ is cluster-positive.
Lemma 4.8.
Let $S=(p,q,r)\in\mathcal{S}^{+}$ be cluster-positive. Then, we have $p,q,r\geq 2$.
Lemma 4.9.
Let $S\in\mathcal{S}^{+}$ be cluster-positive. Then, we have $C(S)\leq 4$.
5. Proof of Lemma 4.7
Here, we show Lemma 4.7. The essential part is as follows.
Lemma 5.1.
Let $S=(p,q,r)\in\mathcal{S}^{+}$ satisfy $q,r\geq 2$ and $C(S)\leq 4$. Then, we have $p\geq 2$.
Proof.
We have
(31) $C(S)=p^{2}-pqr+q^{2}+r^{2}=\left(p-\frac{1}{2}qr\right)^{2}-\frac{1}{4}q^{2}r^{2}+q^{2}+r^{2}.$
Thus, the inequality $C(S)\leq 4$ turns into
(32) $\left(p-\frac{1}{2}qr\right)^{2}\leq\frac{1}{4}q^{2}r^{2}-q^{2}-r^{2}+4=\frac{1}{4}(q^{2}-4)(r^{2}-4).$
Since the last term is non-negative, this inequality implies
(33) $p-\frac{1}{2}qr\geq-\frac{1}{2}\sqrt{(q^{2}-4)(r^{2}-4)}.$
In particular, $f(q,r)=\frac{1}{2}(qr-\sqrt{(q^{2}-4)(r^{2}-4)})$ is a lower bound of $p$. Thus, it suffices to show that $f(q,r)\geq 2$ for any $q,r\geq 2$. The inequality $f(q,r)\geq 2$ is equivalent to
(34) $qr-4\geq\sqrt{(q^{2}-4)(r^{2}-4)}.$
Since both sides are non-negative, it is equivalent to $(qr-4)^{2}\geq(q^{2}-4)(r^{2}-4)$. Then, we have
(35) $(qr-4)^{2}-(q^{2}-4)(r^{2}-4)=4q^{2}+4r^{2}-8qr=4(q-r)^{2}\geq 0.$
It implies $f(q,r)\geq 2$ as we said above. This completes the proof. ∎
By using this lemma, we may show Lemma 4.7 inductively.
Proof of Lemma 4.7.
Let $S=(p,q,r)\in\mathcal{S}^{+}$ satisfy $p,q,r\geq 2$ and $C(S)\leq 4$ and set $S^{\prime}=(p^{\prime},q^{\prime},r^{\prime})=\gamma_{k}(S)$. By definition of $\gamma_{k}$, all entries other than $k$th one do not change. In particular, they are larger or equal to $2$. By Proposition 4.5 (c), $C(S^{\prime})=C(S)\leq 4$ holds. Thus, the $k$th entry of $S^{\prime}$ is also larger or equal to $2$ by Lemma 5.1. By repeating this argument, every entry of elements in $\Gamma(S)$ is larger or equal to $2$ and, in particular, positive. Thus, the claim holds. ∎
6. $1,2$-orbits and proof of Lemma 4.8
This contents was introduced in [FT19, Lem. 3.3].
Definition 6.1.
Let $r$ be an indeterminate. Then, define polynomials $u_{n}(r)$ ($n\in\mathbb{Z}_{\geq-2}$) recursively as follows:
(36) $u_{-2}(r)$ $=-1,\ u_{-1}(r)=0,$
$u_{n+1}(r)$ $=ru_{n}(r)-u_{n-1}(r)\quad(n\geq-1).$
Note that $u_{0}(r)=1$ and $u_{1}(r)=r$. These polynomials are closely related to the Chebyshev polynomials $U_{n}(r)$ of the second kind, which are obtained as the following recursions:
(37) $U_{-2}(r)$ $=-1,\ U_{-1}(r)=0,$
$U_{n+1}(r)$ $=2rU_{n}(r)-U_{n-1}(r)\quad(n\geq-1).$
Note that $U_{0}(r)=1$ and $U_{1}(r)=2r$. In particular, we have
(38) $u_{n}(r)=U_{n}\left(\frac{r}{2}\right).$
For any triplet $S=(p,q,r)\in\mathcal{S}$ and integer $n\geq-1$, set $f_{n}(S)=qu_{n}(r)-pu_{n-1}(r)$. Then, we have $f_{-1}(S)=p$ and $f_{0}(S)=q$. Moreover, $f_{n}(S)$ obeys the following recursion:
(39) $f_{n+1}(S)=rf_{n}(S)-f_{n-1}(S).$
Definition 6.2.
Let $S\in\mathcal{S}^{+}$. We define a collection $\Gamma_{1,2}(S)=\{\gamma_{t}(S)\}_{t}$ indexed by a sequence $t=(t_{1},\dots,t_{l})$ of $1,2$ such that $t_{i}\neq t_{i+1}$ for any $i=1,\dots,l-1$.
Every element of $\Gamma_{1,2}(S)$ may be expressed as follows:
Proposition 6.3 ([FT19, Lem. 3.3]).
Fix $S=(p,q,r)$, and let $p_{n}$ and $q_{n}$ ($n\geq 0$) be
(40) $(p_{n},q_{n},r)=\begin{cases}(\gamma_{2}\gamma_{1})^{\frac{n}{2}}(S)&\textup{$n$ is even},\\ \gamma_{1}(\gamma_{2}\gamma_{1})^{\frac{n-1}{2}}(S)&\textup{$n$ is odd}.\end{cases}$
Then, for any $n\in\mathbb{Z}_{\geq 0}$, we have
(41) $p_{n}=\begin{cases}f_{n-1}(S)&\textup{$n$ is even},\\ f_{n}(S)&\textup{$n$ is odd},\end{cases}\quad q_{n}=\begin{cases}f_{n}(S)&\textup{$n$ is even},\\ f_{n-1}(S)&\textup{$n$ is odd}.\end{cases}$
We may check it by the induction on $n$. By using this property, entries of $\Gamma_{1,2}(S)$ are closely related to the Chebychev polynomial. The following properties are well known for the Chebyshev polynomials.
Proposition 6.4 (e.g. [HM03, (1.4), (1.33b)]).
For any $n\geq 0$ and $\theta\in\mathbb{R}$, we have $\sin(n+1)\theta=\sin\theta U_{n}(\cos\theta)$ and $\sinh(n+1)\theta=\sinh\theta U_{n}(\cosh\theta)$. Moreover, we have $U_{n}(1)=n+1$.
We give some fundamental properties of $f_{n}(S)$.
Proposition 6.5 ([FT19, Lem. 3.3]).
Let $S=(p,q,r)\in\mathcal{S}^{+}$. If $r<2$, then there is a negative entry in $\Gamma_{1,2}(S)$.
Proof.
For the reader’s convenience, we give a proof following [FT19]. When $20<r<$ 2, we may express $r=2\cos\theta$ for some $\theta\in\mathbb{R}$ with $0<\theta<\frac{\pi}{2}$. By Proposition 6.3, $f_{n}(S)=qu_{n}(r)-pu_{n-1}(r)$ is a entry of an element in $\Gamma(S)$. We show that there is $n\in\mathbb{Z}_{\geq 0}$ such that $f_{n}(S)<0$. Note that $u_{n}(r)=U_{n}(\cos\theta)=\frac{\sin(n+1)\theta}{\sin\theta}$ by (38) and Proposition 6.4. By using this relation, the inequality $f_{n}(S)<0$ is equivalent to $q\sin(n+1)\theta<p\sin n\theta$. Here we use $\sin\theta>0$ since $0<\theta<\frac{\pi}{2}$. We can find $n\geq 1$ such that $\sin(n+1)\theta\leq 0$ while every $k=1,\dots,n$ satifies $\sin k\theta>0$. This $n$ satisfies $q\sin(n+1)\theta<p\sin n\theta$. ∎
By using this property, Lemma 4.8 is immediately shown as follows.
Proof of Lemma 4.8.
More precisely, we show the contraposition of Lemma 4.8. Suppose that $S=(p,q,r)\in\mathcal{S}^{+}$ satisfies $r<2$. Then, by Proposition 6.5, there is a negative entry in $\Gamma_{1,2}(S)\subset\Gamma(S)$. Thus, $S$ is not cluster-positive. ∎
Proposition 6.6.
Let $S=(p,q,r)\in\mathcal{S}^{+}$ with $p,q,r\geq 2$ and set $r=2\cosh\theta$ for some $\theta\in\mathbb{R}_{\geq 0}$. Let $p_{n}$ and $q_{n}$ be in (41). If $p_{n}$ or $q_{n}$ converges to some real number, then we have $qe^{\theta}-p=0$. Moreover,
(a) if $r=2$, then we have $p_{n}=q_{n}=p=q$ for any $n=0,1,2,\dots$.
(b) if $r>2$, then both $p_{n}$ and $q_{n}$ converge to 00.
Proof.
Suppose that $p_{n}$ converges. This is equivalent that $f_{2k+1}$ ($k\in\mathbb{Z}_{>0}$) converges. Let $n=2k+1$. When $r=2$, it holds that $\theta=0$. Since $u_{n}(2)=U_{n}(1)=n+1$, we have $f_{n}(S)=q(n+1)-pn=(q-p)n+q$. Thus, if $f_{n}(S)$ converges, we have $q-p=qe^{0}-p=0$. By considering the maps $\gamma_{1}$ and $\gamma_{2}$, we may obtain $p_{n}=q_{n}=p=q$ for any $n\in\mathbb{Z}_{>0}$. If $r>2$, or equivalently, if $\theta>0$, it holds that $\sin n\theta>0$ for any $n\in\mathbb{Z}_{>0}$ and we have $\frac{\sinh(n+1)\theta}{\sinh n\theta}=\frac{e^{(n+1)\theta}-e^{-(n+1)\theta}}{e^{n\theta}-e^{-n\theta}}=\frac{e^{\theta}-e^{-(2n+1)\theta}}{1-e^{-2n\theta}}\to e^{\theta}$ when $n\to\infty$. Since $u_{n}(r)=U_{n}(\cosh\theta)=\frac{\sinh(n+1)\theta}{\sinh n\theta}$, we have
(42) $f_{n}(S)$ $=q\frac{\sinh(n+1)\theta}{\sinh\theta}-p\frac{\sinh n\theta}{\sinh\theta}$
$=\frac{\sinh n\theta}{\sinh\theta}\left(q\frac{\sinh(n+1)\theta}{\sinh n\theta}-p\right).$
Unless $qe^{\theta}-p\neq 0$, it does not converge since $\sinh(2n+1)\theta\to\infty$ when $n\to\infty$. When $p=qe^{\theta}$, we have
(43) $f_{n}(S)$ $=q\frac{\sinh(n+1)\theta}{\sinh\theta}-qe^{\theta}\frac{\sinh n\theta}{\sinh\theta}$
$=\frac{q}{\sinh\theta}\left(\sinh(n+1)\theta-e^{\theta}\sinh\theta\right)$
$=\frac{q}{2\sinh\theta}\left(-e^{-(n+1)\theta}+e^{(-n+1)\theta}\right)\to 0.$
[…]
Later, we will show $C(S)\leq 4$ for any cluster-positive element $S$. It was shown in [FT19] by using the geometrical aspects of mutations. Here, we give a proof which is more similar to that in [BBH11], althoug we need some modifications for the real case.
In the previous section, for any $S\in\mathcal{S}^{+}$, $\Gamma(S)=\{\gamma_{t}(S)\}_{t}$ was defined as a collection indexed by $t$. However, in this section, we view $\Gamma(S)=\{\gamma_{t}(S)\mid t\}$ as a set.
Let $\leq$ be a partial order on $\mathcal{S}$ such that every corresponding entry satisfies this inequality on $\mathbb{R}$.
Definition 7.1.
[…]
  • (M1)
    For any $i=1,2,3$, it holds that $S\leq\gamma_{i}(S)$.
  • (M2)
    For presicely two indices $i$ in $\{1,2,3\}$, it holds that $S\leq\gamma_{i}(S)$.
  • (M3)
    Otherwise, namely, at most one index $i$ in {1,2,3} satisfies $S\leq\gamma_{i}(S)$.
Acyclic matrices are closely related to (M3) as follows.
Lemma 7.2.
Let $S=(p,q,r)\in\mathcal{S}$. If $p,q>0$ and $r\leq 0$, then $S$ satisfies (M3).
Proof.
We have
(44) $\gamma_{1}(S)=(qr-p,q,r),\ \gamma_{2}(S)=(p,pr-q,r).$
Then, since $qr-p<0<p$ and $pr-q<0<q$, this $S$ satisfies $S\nleq\gamma_{1}(S)$ and $S\nleq\gamma_{2}(S)$. It implies that $S$ satisfies (M3). ∎
[…]
Proposition 7.3.
Let $S=(p,q,r)\in\mathcal{S}^{+}$ satisfy $p\geq q\geq r$.
(1) A triplet $S$ satisfies (M1) if and only if $S$ satisfies $S\leq\gamma_{1}(S)$, or equivalently, $qr\geq 2p$.
(2) A triplet $S$ satisfies (M2) if and only if $2p^{2}>pqr\geq 2q^{2}$. In this case, $S$ satisfies $S\nleq\gamma_{1}(S)$ and $S\leq\gamma_{2}(S),\gamma_{3}(S)$.
(3) A triplet $S$ satisfies (M3) if and only if $2q>pr$. In this case, $S$ satisfies $S\nleq\gamma_{1}(S),\gamma_{2}(S)$.
Proof.
Note that
$S\leq\gamma_{1}(S)$Leftrightarrow $qr-p\geq p$Leftrightarrow $qr\geq 2p$Leftrightarrow $pqr\geq 2p^{2}$,
$S\leq\gamma_{2}(S)$Leftrightarrow $pqr\geq 2q^{2}$,
$S\leq\gamma_{3}(S)$Leftrightarrow $pqr\geq 2r^{2}$.
Under the assumption $p\geq q\geq r>0$, whether $S$ satisfies (M1), (M2), or (M3) is determined by where $pqr$ is in $2p^{2}\geq 2q^{2}\geq 2r^{2}$. By considering each condition, we may show Proposition 7.3
Following a similar argument as in [BBH11, Lem. 2.1], we will prove some fundamental properties of these classes in $\mathcal{S}^{+}$.
Lemma 7.4 (see [BBH11, Lem. 2.1. (c)]).
If $S=(p,q,r)\in\mathcal{S}^{+}$ satisfies (M3), then we have $\min(p,q,r)<2$ and $S$ is not cluster-positive.
Proof.
Without loss of generality, we may assume $p\geq q\geq r>0$. By Proposition 7.3 (3), we have $2q>pr$ and $2p>qr$. By multiplying the corresponding sides, we have
(45) $4pq>pqr^{2}$
and it implies $4>r^{2}$, namely, $r<2$. By Proposition 6.5, we may verify that $S$ is not cluster-positive. ∎
Lemma 7.5 (see [BBH11, Lem 2.1 (a)]).
Let $S=(p,q,r)\in\mathcal{S}^{+}$ satisfy (M1). Then, the following statements hold.
(a) We have three inequarities $p,q,r\geq 2$.
(b) Every element in $\Gamma(S)$ satisfies (M2) except for $S$. Here, we see $\Gamma(S)=\{\gamma_{t}(S)\mid t\}$ as a set.
(c) For any sequence
(46) $S=S_{0}\overset{\gamma_{t_{1}}}{\mapsto}S_{1}\overset{\gamma_{t_{2}}}{\mapsto}S_{2}\overset{\gamma_{t_{3}}}{\mapsto}\cdots\overset{\gamma_{t_{k}}}{\mapsto}S_{k}=S^{\prime}$
which satisfies $t_{i-1}\neq t_{i}$ ($i=1,2,\dots,k$), then we have $S_{i-1}\leq S_{i}$. In particular, for any $S^{\prime}\in\Gamma(S)$, there is a sequence (46) with $S_{i-1}\leq S_{i}$, and $S_{i-1}\neq S_{i}$ for any $i=1,2,\dots,k$.
(d) This $S$ is the smallest element in $\Gamma(S)$ in terms of the partial order $\leq$. In particular, $S$ is cluster-positive.
Proof.
First, we show (a). By Proposition 7.3 (1), we have
(47) $pqr\geq 2p^{2},2q^{2},2r^{2}$
For example, by multiplying both sides of two inequatries $pqr\geq 2p^{2}$ and $pqr\geq 2q^{2}$, we have $(pqr)^{2}\geq 4p^{2}q^{2}$, and it implies $r^{2}\geq 4$ and $r\geq 2$. Similarly, we have $p,q\geq 2$.
We show (b) and (c) by the induction on the distance $d$ from $S$ to $S^{\prime}$. When $d=0$, namely, $S=S^{\prime}$, it is obvious. Suppose that the claim holds for some $d$ and let the distance form $S$ to $S^{\prime}$ be $d+1$. Then, we can find $t_{d+1}=1,2,3$ and $S_{d}=\gamma_{t_{d+1}}(S^{\prime})$ such that the distance from $S$ to $S_{d}$ is $d$. By the assumption, we obtain a sequence $S,S_{1},\dots,S_{d}$ satisfying (c) and $S_{i}$ satisfies (M2) for any $i=1,2,\dots,d$. Let $d\geq 1$. Since $S_{d-1}\leq S_{d}=\gamma_{t_{d}}(S_{d-1})$ and $S_{d-1}\neq S_{d}$, we have $S_{d}\nleq S_{d-1}=\gamma_{t_{d}}(S_{d})$. Thus, since $S_{d}$ satisfies (M2) and $t_{d}\neq t_{d+1}$, we have $S_{d}\leq S_{d+1}$. This inequality also holds for $d=0$ since $S=S_{0}$ satisfies (M1). If $S_{d}=S^{\prime}$, the sequence $S,S_{1},\dots,S_{d}=S^{\prime}$ satisfies (c). If $S_{d}\neq S^{\prime}$, then the sequence $S,S_{1},\dots,S_{d},S^{\prime}$ satisfies (c). Thus, (c) holds. By considering this sequence, $S^{\prime}$ does not satisfy (M1) since $S^{\prime}\nleq\gamma_{t_{d+1}}(S^{\prime})$. Moreover, $S^{\prime}$ does not satisfy (M3). This is because, if $S^{\prime}=(p^{\prime},q^{\prime},r^{\prime})$ satisfies (M3), then we have $\min(p^{\prime},q^{\prime},r^{\prime})<2$ by Lemma 7.4, but it contradicts $(2,2,2)\leq S\leq S^{\prime}$ by (a) and (c).
Last, we show (d). By (c), we have $S\leq S^{\prime}$ for any $S^{\prime}\in\Gamma(S)$. Thus, $S$ is the smallest element of $\Gamma(S)$, and it is cluster-positive. ∎
By combining these results, we can give two classes of cluster-positive elements.
Definition 7.6.
For any cluster-positive element $S\in\mathcal{S}^{+}$, we define the following conditions.
  • (A)
    $\Gamma(S)$ has a unique element satisfying (M1).
  • (B)
    All elements of $\Gamma(S)$ satisfy (M2).
By Lemma 7.4 and Lemma 7.5, every cluster-positive element satisfies either (A) or (B). For the case of (B), there is a sequence
(48) $S_{0}=S,S_{1}=\gamma_{t_{1}}(S_{0}),S_{2}=\gamma_{t_{2}}(S_{1}),\dots\quad(t_{i}\neq t_{i+1}\ \textup{for any $i$})$
such that $S_{i}$ satisfies (M2) and $S_{i-1}\nleq S_{i}$ for any $i$. Moreover, this sequence is uniquely determined by $S\in\mathcal{S}^{+}$.
Now we identify $\mathcal{S}^{+}$ as >$\mathbb{R}^{3}_{>0}$, and we introduce a topology on $\mathcal{S}^{+}$ by this correspondence. Then, the limit of this sequence also satisfies (M1).
Lemma 7.7.
Let $S\in\mathcal{S}^{+}$ satisfy the condition (B) and let $\{S_{i}\}$ be a sequence of (48). Then, we have the following statements.
(a) The sequence $\{S_{i}\}$ converges to $(2,2,2)$.
(b) We have $C(S)=4$.
The existence of the limit was shown in [BBH11, Thm. 5.1], but we can show that this limit is $(2,2,2)$.
Proof.
(a) First, we show that $S_{i}$ converges to some $L(S)=(a,b,c)$ with $a,b,c\geq 2$. Let $S_{i}=(p_{i},q_{i},r_{i})$. Then, since $S_{i}$ is cluster-positive, we have $p_{i},q_{i},r_{i}\geq 2$. Moreover, these sequences $\{p_{i}\}$, $\{q_{i}\}$, and $\{r_{i}\}$ are monotonically decreasing. Thus, they converges to some real numbers $a$, $b$, and $c$, respectively. Thus, $S_{i}$ converges to $L(S)=(a,b,c)$. Moreover, since $p_{i},q_{i},r_{i}\geq 2$, we have $a,b,c\geq 2$. Next, we show $a=b=c=2$. Without loss of generality, we may assume $a\geq b\geq c$. Since monotonically decreasing sequences $p_{i}$, $q_{i}$, and $r_{i}$ converges to $a$, $b$, and $c$, respectively, for any $\varepsilon>0$, there is $N\in\mathbb{Z}_{>0}$ such that $a\leq p_{i}<a+\varepsilon$, $b\leq q_{i}<b+\varepsilon$, and $c\leq r_{i}<c+\varepsilon$ for any $i\geq N$. Since $S_{i}$ satisfies (M2), there is $i\geq N$ such that $p_{i}>q_{i}\geq r_{i}$. By Proposition 7.3 (2), we have $q_{i}r_{i}<2p_{i}$. Since $p_{i}<a+\varepsilon$ and $bc\leq q_{i}r_{i}$, we have $bc<2(a+\varepsilon)$. This inequality holds for arbtrary $\varepsilon>0$. By considering the limit $\varepsilon\to+0$, we have $bc\leq 2a$. Next, consider $S_{i+1}$. Since $p_{i}>q_{i}\geq r_{i}$ and $S_{i}\nleq S_{i+1}$, we need to choose $S_{i+1}=\gamma_{1}(S_{i})$. In order $S_{i+1}$ to satisfy (M2), we need $q_{i+1}>p_{i+1},r_{i+1}$. Thus, by doing a similar argument for $S_{i}$, we have $ac\leq 2b$. By the two inequalities $bc\leq 2a$ and $ac\leq 2b$, we have $c^{2}\leq 4$ and, by $c\geq 2$, we have $c=2$. This implies that $a=b$. Now, we assume $a=b>2=c$. Then, we have $p_{i},q_{i}>2=r_{i}$ for any $i$, and we should choose the sequence of indices as $(1,2,1,2,\dots)$ or $(2,1,2,1,\dots)$. By Proposition 6.6, we have $p_{i}=q_{i}=a$ because $S_{i}$ converges to some nonzero number $a$. However, $S_{i}=(a,a,2)$ satisfies (M1), and this contradicts the fact that $S_{i}$ satisfies (M2). Thus, we have $a=b=c=2$.
(b) Note that $C(S)=C(S_{i})=p_{i}^{2}+q_{i}^{2}+r_{i}^{2}-p_{i}q_{i}r_{i}$ does not depend on $i$. By considering the limit of $i\to\infty$, we have $C(S)=2^{2}+2^{2}+2^{2}-2\cdot 2\cdot 2=4=C(L(S))$. ∎
Based on this result, we define the following map.
Definition 7.8.
For any cluster-positive element $S\in\mathcal{S}^{+}$, let $L(S)\in\mathcal{S}^{+}$ as follows:
  • If $S$ satisfies (A), let $L(S)$ be the element of $\Gamma(S)$ satisfying (M1). (By Lemma 7.5, this is uniquely determined by $S$.)
  • If $S$ satisfies (B), define $L(S)=(2,2,2)$.
This $L(S)$ is like an infimum of $\Gamma(S)$ for any cluster-positive element $S\in\mathcal{S}^{+}$. Note that, for any $S$ satisfying (A), then $L(S)$ is the smallest element in $\Gamma(S)$ (Lemma 7.5) and satisfies (M1). We summarize propeties of $L(S)$, which was shown in Lemma 7.5 and Lemma 7.7.
Proposition 7.9.
Let $S\in\mathcal{S}^{+}$ be cluster-positive. Then, we have the following properties.
(a) We have $C(S)=C(L(S))$.
(b) $L(S)$ satisfies (M1).
(c) For any $S^{\prime}\in\Gamma(S)$, we have $L(S)\leq S^{\prime}$.
8. Proof of Lemma 4.9
Let $S\in\mathcal{S}^{+}$ be cluster-positive. Then, by Proposition 7.9, there is $L(S)$ satisfying (M1) such that $C(S)=C(L(S))$. Thus, to show Lemma 4.9, it suffices to show the following statement.
Lemma 8.1.
If $S\in\mathcal{S}^{+}$ satisfies (M1), we have $C(S)\leq 4$.
For later, we show the following lemma at the same time.
Lemma 8.2 (cf. [BBH11, Lem. 3.3]).
Let $S=(p,q,r)\in\mathcal{S}^{+}$ satisfy (M1), and we assume $p\geq q\geq r$.
(1) We have $C(S)\leq C\left(r,r,r\right)$.
(2) The equality $C(S)=4$ holds if and only if $S=(p,p,2)$ for some $p\geq 2$.
In [BBH11], they showed it for any integer triplet $p,q,r\in\mathbb{Z}$. This result is a generalization for the real case. The following proof is similar but slightly different to the one of [BBH11, Lem. 3.3]. In particular, the proof is different when we consider $32<r<$ 3.
Proof.
Since $S$ satisfies (M1), by Lemma 7.5 (a) and Proposition 7.3 (1), we have $\frac{1}{2}qr\geq p\geq q\geq r\geq 2$. Now, we fix $r\geq 2$ and we see $C(S)$ as a function of $(p,q)\in\mathbb{R}^{2}$ satisfying the above inequalities. The region where $(p,q)$ satisfies these inequalities is illustrated in Figure 3. Set $f(p,q)=p^{2}+q^{2}+r^{2}-pqr$.
$p$$q$$qr=2p$$p=q$$r$$q$$(p_{1},q)$$(p,q)$$(p_{2},q)$
Figure 3. The region of $(p,q)$
First, we show Lemma 8.2 (1). We fix $q\geq r$ and we view $f(p,q)$ as a function of $p$. Set $p_{1}=q$ and $p_{2}=\frac{1}{2}qr$. In this case, $p$ satisfies $p_{1}\leq p\leq p_{2}$ as in Figure 3. Then, since $f(p,q)$ is a quadratic function with the positive leading coefficient $1$, we have $f(p,q)\leq\max(f(p_{1},q),f(p_{2},q))$. We can show $f(p_{1},q)\geq f(p_{2},q)$ as follows:
(49) $f(p_{1},q)-f(p_{2},q)$ $=(p_{1})^{2}-(p_{2})^{2}-(p_{1}qr-p_{2}qr)$
$=(p_{1}-p_{2})(p_{1}+p_{2})-qr(p_{1}-p_{2})$
$=(p_{1}-p_{2})(p_{1}-qr+p_{2})\geq 0.$
The last inequality may be obtained by $p_{1}-p_{2}\leq 0$ and $p_{1}-qr+p_{2}=q-qr+\frac{1}{2}qr=-\frac{1}{2}q(r-2)\leq 0$. Thus, we have $f(p,q)\leq f(p_{1},q)=f(q,q)$. Moreover, since $f(q,q)=2q^{2}+r^{2}-q^{2}r=q^{2}(2-r)+r^{2}$ is a quadratic function of $q$ with the negative coefficient $02-r<$ 0 or a constant $r^{2}$, we have $f(q,q)\leq f(r,r)=C(r,r,r)$. (We used $q\geq r>0$.) Thus, we have $C(S)\leq C(r,r,r)$.
Next, we show Lemma 8.1. By Lemma 8.2 (1), it suffices to show that $C(r,r,r)=3r^{2}-r^{3}\leq 4$. It may be shown as follows:
(50) $4-(3r^{2}-r^{3})=(r+1)(r-2)^{2}\geq 0.\quad(r\geq 2)$
Last, we show Lemma 8.2 (2). Since $C(S)\leq C(r,r,r)\leq 4$, we need $C(r,r,r)=4$ for $C(S)=4$. Moreover, $C(r,r,r)=4$ turns into the equality $(r+1)(r-2)^{2}=0$. By $r\geq 2$, we need $r=2$ to be $C(S)=4$. In this case, we have
(51) $C(S)=p^{2}+q^{2}+4-2pqr=(p-q)^{2}+4.$
Thus, $C(S)=4$ is equivalent to $p=q\geq r=2$. ∎
9. The case (B)
Recall that there are two classes (A) and (B) of cluster-positive elements. (See Definition 7.6.) By considering Lemma 7.7 and Lemma 8.2, we have the following theorem.
Theorem 9.1.
Let $S\in\mathcal{S}^{+}$ be cluster-positive.
(a) This element $S$ satisfies (A) if and only if $C(S)<4$ or $S$ belongs to $\Gamma(p,p,2)$, $\Gamma(p,2,p)$, or $\Gamma(2,p,p)$ for some $p\geq 2$.
(b) This element $S\in\mathcal{S}^{+}$ satisfies (B) if and only if $C(S)=4$ and $S$ does not belong to $\Gamma(p,p,2)$, $\Gamma(p,2,p)$, and $\Gamma(2,p,p)$ for any $p\geq 2$.
Proof.
By Lemma 7.7, if a cluster-positive element $S$ satisfies (B), then we have $C(S)=4$. Moreover, cluster-positive elements $S\in\mathcal{S}^{+}$ satisfying $C(S)=4$ and (A) should belong to $\Gamma(p,p,2)$, $\Gamma(p,2,p)$, or $\Gamma(2,p,p)$ by Lemma 8.2 (2). Thus, the claim holds. ∎
An element $S=(\frac{29}{2},\frac{19}{4},\frac{7}{2})$ would satisfy (B). In fact, I have already checked that there is a decreasing sequence $S_{0},S_{1},\dots,S_{50}$ by using a computer. However, I cannot show that this $S$ satisfies (B). (We may check $C(S)=4$, but it would be difficult to show $S\notin\Gamma(p,p,2)\cup\Gamma(p,2,p)\cup\Gamma(2,p,p)$ for any $p\geq 2$.)
In [FT19, Rem. 4.7], another identification between (A) and (B) is shown.
10. Integer skew-symmetrizable case
Here, we consider integer skew-symmetrizable matrices and their orbits by $\gamma_{k}$. For this, we consider $\Gamma$ as a subgroup of $\mathrm{Aut}(\mathcal{M})$ generated by $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$.
Recall that $\mathcal{M}$ is defined by the set of all $M=\left(\begin{smallmatrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{smallmatrix}\right)$ corresponding to a skew-symmetrizable matrix $B\in\mathrm{M}_{3}(\mathbb{R})$ as in Defnition 4.2. In this section, we consider the following subsets.
Definition 10.1.
Let $\mathcal{M}_{\mathbb{Z}}$ (resp. $\mathcal{M}^{+}_{\mathbb{Z}}$) be the set of all $M\in\mathcal{M}$ (resp. $M\in\mathcal{M}^{+}$) whose entries are integral. Moreover, we write the set of all cluster-positive elements in $\mathcal{M}_{\mathbb{Z}}$ by $\mathcal{M}_{\mathbb{Z}}^{\mathrm{CP}}$.
For the sake of simplicity, we introduce an action of the symmetric group $\mathfrak{S}_{3}$ on $\mathcal{M}$ as
(52) $\sigma\left(\begin{matrix}a_{1}&a_{2}&a_{3}\\ a^{\prime}_{1}&a^{\prime}_{2}&a^{\prime}_{3}\end{matrix}\right)=\left(\begin{matrix}a_{\sigma^{-1}(1)}&a_{\sigma^{-1}(2)}&a_{\sigma^{-1}(3)}\\ a^{\prime}_{\sigma^{-1}(1)}&a^{\prime}_{\sigma^{-1}(2)}&a^{\prime}_{\sigma^{-1}(3)}\end{matrix}\right),$
where $\sigma\in\mathfrak{S}_{3}$. We introduce another action $t:\mathcal{M}\to\mathcal{M}$ by $t\left(\begin{smallmatrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{smallmatrix}\right)=\left(\begin{smallmatrix}x^{\prime}&y^{\prime}&z^{\prime}\\ x&y&z\end{smallmatrix}\right)$. Then, every element obtained by applying these actions for $M\in\mathcal{M}$ is called a permutation of $M$. Similarly, we define permutations of $S\in\mathcal{S}$. (For $\mathcal{S}$, we set $t=\mathrm{id}$.)
10.1. Fundamental domain
Recall that $\mathcal{S}$ is defined by the set of all $S=(p,q,r)\in\mathbb{R}^{3}$. Then, there is a map $\mathrm{Sk}:\mathcal{M}\to\mathcal{S}$ with $\mathrm{Sk}\left(\begin{smallmatrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{smallmatrix}\right)=(\varepsilon_{x}\sqrt{xx^{\prime}},\varepsilon_{y}\sqrt{yy^{\prime}},\varepsilon_{z}\sqrt{zz^{\prime}})$, where $\varepsilon_{a}\in\{0,\pm 1\}$ is the sign of $a\in\mathbb{R}$. By using this map, we define $\widehat{\mathcal{S}}=\mathrm{Sk}({\mathcal{M}}_{\mathbb{Z}})$, $\widehat{\mathcal{S}}^{+}=\mathrm{Sk}({\mathcal{M}}_{\mathbb{Z}}^{+})$, and $\widehat{\mathcal{S}}^{\mathrm{CP}}=\mathrm{Sk}({\mathcal{M}}_{\mathbb{Z}}^{\mathrm{CP}})$.
First, we give an expression of $\widehat{\mathcal{S}}$.
Proposition 10.2.
We have
(53) $\widehat{\mathcal{S}}$ =$=\{(p,q,r)\in\mathcal{S}\mid p^{2},q^{2},r^{2},pqr\in\mathbb{Z}\}.$
Proof.
The inclusion $\widehat{\mathcal{S}}\subset\{(p,q,r)\mid p^{2},q^{2},r^{2},pqr\in\mathbb{Z}\}$ is immediately shown by the definition of $\mathrm{Sk}(M)=(p,q,r)$. Note that $pqr\in\mathbb{Z}$ follows from (10). Let $S=(p,q,r)\in\widehat{\mathcal{S}}$. We find $M\in\mathcal{M}_{\mathbb{Z}}$ with $\mathrm{Sk}(M)=S$. If at least one entry of $S$ is zero, for example $r=0$, take $M=\left(\begin{smallmatrix}p^{2}&1&0\\ 1&q^{2}&0\end{smallmatrix}\right)\in\mathcal{M}_{\mathbb{Z}}$. Assume $p,q,r\neq 0$. Since $p^{2},q^{2},r^{2}\in\mathbb{Z}$, we may express $p=l\sqrt{a}$, $q=m\sqrt{b}$, and $r=n\sqrt{c}$, where $\0l,m,n\in\mathbb{Z}\backslash\{0\}$ and $a,b,c\in\mathbb{Z}_{>0}$ do not have any square factors. Since $pqr=lmn\sqrt{abc}\in\mathbb{Z}$ and $\0lmn\in\mathbb{Z}\backslash\{0\}$, we need $\sqrt{abc}\in\mathbb{Z}$. Let $p_{k}$ be the $k$th prime number, and consider factorizations $a=\prod p_{k}^{\varepsilon_{k}^{a}}$, $b=\prod p_{k}^{\varepsilon_{k}^{b}}$, and $c=\prod p_{k}^{\varepsilon_{k}^{c}}$. Since $a$, $b$, and $c$ do not have square factors, every exponent of $p_{k}$ should be 00 or $1$. Moreover, since $\sqrt{abc}=\sqrt{\prod p_{k}^{\varepsilon_{k}^{a}+\varepsilon_{k}^{b}+\varepsilon_{k}^{c}}}$ is an integer, we need $\varepsilon_{k}^{a}+\varepsilon_{k}^{b}+\varepsilon_{k}^{c}\in 2\mathbb{Z}$ for any $k$. In particular, $(\varepsilon_{k}^{a},\varepsilon_{k}^{b},\varepsilon_{k}^{c})$ is any of $(0,0,0)$, $(1,1,0)$, $(0,1,1)$, or $(1,0,1)$. Let
(54) $u=\prod_{(\varepsilon_{k}^{a},\varepsilon_{k}^{b})=(1,1)}p_{k},\ v=\prod_{(\varepsilon_{k}^{b},\varepsilon_{k}^{c})=(1,1)}p_{k},\ w=\prod_{(\varepsilon_{k}^{c},\varepsilon_{k}^{a})=(1,1)}p_{k}.$
Then, we have $a=wu$, $b=uv$, and $c=vw$. By setting
(55) $M=\left(\begin{matrix}lu&mv&nw\\ lw&mu&nv\end{matrix}\right)\in\mathcal{M}_{\mathbb{Z}},$
we have $\mathrm{Sk}(M)=S$. ∎
As Definition 7.6, every cluster-positive element $S\in\mathcal{S}$ satisfies the condition (A) or (B). However, when we consider $S\in\widehat{\mathcal{S}}$, the case (B) does not occur as follows. (This proposition is a generalization of [BBH11, Lem. 2.1 (b)] for the skew-symmetrizable case.)
Proposition 10.3 (cf. [BBH11, Lem. 2.1 (b)]).
If $S\in\widehat{\mathcal{S}}^{+}$ is cluster-positive, then $S$ satisfies (A), which is defined in Definition 7.6.
Proof.
Suppose that there exists a monotonically decreasing sequence (48) and set $S_{i}=(p_{i},q_{i},r_{i})\in\widehat{\mathcal{S}}^{+}$. Then, their product $p_{i}q_{i}r_{i}$ is strictly monotonically decreasing. Since $p_{i}q_{i}r_{i}\in\mathbb{Z}$, it implies that $p_{i}q_{i}r_{i}$ must be negative if $i$ is enough large. Thus, $S$ must not be cluster-positive. ∎
Moreover, by Lemma 7.5, an element in $\Gamma(S)$ satisfying (M1) is unique for any $S\in\widehat{\mathcal{S}}^{\mathrm{CP}}$. In particular, the set
[…]
is a fundamental domain for the action of $\Gamma$ on $\widehat{\mathcal{S}}^{\mathrm{CP}}$ [BBH11, Thm. 5.1]. We generalize this result for the skew-symmetrizable case.
For this, we introduce a partial order $\leq$ on $\mathcal{M}$ as each corresponding entry satisfies this order on $\mathbb{R}$. Then, this order is compatible for the order on $\mathcal{S}$ and the map $\mathrm{Sk}$.
Lemma 10.4.
Let $M=\left(\begin{smallmatrix}a_{1}&a_{2}&a_{3}\\ a^{\prime}_{1}&a^{\prime}_{2}&a^{\prime}_{3}\end{smallmatrix}\right)\in{\mathcal{M}}^{+}$. Then, for any $i=1,2,3$, the following three inequalities are equivalent.
  • (a)
    $M\leq\gamma_{i}(M)$ on $\mathcal{M}$.
  • (b)
    $\mathrm{Sk}(M)\leq\gamma_{i}(\mathrm{Sk}(M))$ on $\mathcal{S}$.
  • (c)
    $a_{1}a_{2}a_{3}\geq 2a_{i}a^{\prime}_{i}$.
Proof.
We show the claim for the case of $i=1$. The condition (a) is equivalent to the condition that both inequalities $a^{\prime}_{2}a^{\prime}_{3}-a_{1}\geq a_{1}$ and $a_{2}a_{3}-a^{\prime}_{1}\geq a^{\prime}_{1}$ hold. These two inequalities are equivalent to $a_{1}a_{2}a_{3}\geq 2a_{1}a^{\prime}_{1}$ since $a_{1}a_{2}a_{3}=a^{\prime}_{1}a^{\prime}_{2}a^{\prime}_{3}$ and $a_{1},a^{\prime}_{1}>0$. Thus, (a) and (c) are equivalent. Similarly, we can show that (b) and (c) are equivalent. ∎
Definition 10.5.
For any $M=\left(\begin{smallmatrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{smallmatrix}\right)\in\mathcal{M}^{+}$, we say that $M$ satisfies (M1) if $\mathrm{Sk}(M)$ satisfies (M1) in terms of Definition 7.1.
By Proposition 7.3 and Lemma 10.4, $M=\left(\begin{smallmatrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{smallmatrix}\right)\in\mathcal{M}^{+}$ satisfies (M1) if and only if $xyz\geq 2xx^{\prime},2yy^{\prime},2zz^{\prime}$.
Now, we can generalize Lemma 7.5 (c) for $\mathcal{M}$.
Lemma 10.6.
Let $M\in\mathcal{M}_{\mathbb{Z}}$ satisfy (M1). Then, for any $M^{\prime}\in\Gamma(M)$, there is a sequence
(57) $M=M_{0}\overset{\gamma_{t_{1}}}{\mapsto}M_{1}\overset{\gamma_{t_{2}}}{\mapsto}\cdots\overset{\gamma_{t_{k}}}{\mapsto}M_{k}=M^{\prime}$
satisfying $M_{i-1}\leq M_{i}$ and $M_{i-1}\neq M_{i}$ for any $i=1,2,\dots,k$. In particular, we have $M\leq M^{\prime}$ for any $M^{\prime}\in\Gamma(M)$.
Proof.
Since $M^{\prime}\in\Gamma(M)$, there is a sequence (57) with $M_{i-1}\neq M_{i}$ and $t_{i-1}\neq t_{i}$ for any $i=1,2,\dots,k$. We show $M_{i-1}\leq M_{i}$. Consider a sequence defined by $S_{i}=\mathrm{Sk}(M_{i})$. Then, by Lemma 7.5 (c), we have $S_{i-1}\leq S_{i}$, and it implies that $M_{i-1}\leq M_{i}$ by Lemma 10.4. ∎
Theorem 10.7.
A set
[…]
=$=\left\{\left(\begin{matrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{matrix}\right)\in{\mathcal{M}}_{\mathbb{Z}}^{+}\mid xyz\geq 2xx^{\prime},2yy^{\prime},2zz^{\prime}\right\}$
is a fundamental domain for the action of $\Gamma$ on $\mathcal{M}_{\mathbb{Z}}^{\mathrm{CP}}$. (Namely, $F\subset\mathcal{M}_{\mathbb{Z}}^{\mathrm{CP}}$ holds and, for any $M\in{\mathcal{M}}_{\mathbb{Z}}^{\mathrm{CP}}$, the set $\Gamma(M)\cap F$ is a singleton.) Moreover, the element of $\Gamma(M)\cap F$ is the smallest element in $\Gamma(M)$.
Proof.
For any $M\in F$, $\mathrm{Sk}(M)$ is cluster-positive since it satisfies (M1). Thus, $M$ is also cluster-positive, namely, the inclusion $F\subset\mathcal{M}_{\mathbb{Z}}^{\mathrm{CP}}$ holds.
Let $M\in\mathcal{M}_{\mathbb{Z}}^{\mathrm{CP}}$. We show $\Gamma(M)\cap F\neq\emptyset$. By Proposition 10.3, there is $S=\gamma_{t}(\mathrm{Sk}(M))\in\Gamma(\mathrm{Sk}(M))$ satisfying (M1) for some finite sequence of indices. Then, since $\gamma_{t}(\mathrm{Sk}(M))=\mathrm{Sk}(\gamma_{t}(M))$, $\gamma_{t}(M)$ satisfies (M1). Thus, we have $\gamma_{t}(M)\in\Gamma(M)\cap F\neq\emptyset$. Next, we show $|\Gamma(M)\cap F|=1$. Without loss of generality, we may suppose that $M$ satisfies (M1). Let $M^{\prime}\in\Gamma(M)\cap F$. Then, by Lemma 10.6, we have $M\leq M^{\prime}$ and $M^{\prime}\leq M$, and it implies that $M=M^{\prime}$. ∎
10.2. Fixed points and finite orbits
Here, we focus on elements $M\in\mathcal{M}^{+}$ with a finite orbit $\Gamma(M)$. In [FST12b, FST12a], they gave a classification of these matrices.
Lemma 10.8.
An element $M=\left(\begin{smallmatrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{smallmatrix}\right)\in\mathcal{M}^{+}$ is a fixed point for the action $\Gamma$ if and only if it holds that
(59) $xx^{\prime}=yy^{\prime}=zz^{\prime}=4.$
Proof.
Let $M=\left(\begin{smallmatrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{smallmatrix}\right)\in\mathcal{M}^{+}$ be a fixed point. Then, by considering $\gamma_{1}(M)=M$, we have $y^{\prime}z^{\prime}=2x$ and $yz=2x^{\prime}$. Since $xyz=x^{\prime}y^{\prime}z^{\prime}$ and $x,x^{\prime}\neq 0$, both equalities are eqivalent to $xyz=x^{\prime}y^{\prime}z^{\prime}=2xx^{\prime}$. Similarly, we have $xyz=x^{\prime}y^{\prime}z^{\prime}=2xx^{\prime}=2yy^{\prime}=2zz^{\prime}$. For example, by multiplying both sides of $xyz=2xx^{\prime}$ and $x^{\prime}y^{\prime}z^{\prime}=2yy^{\prime}$, we have $zz^{\prime}=4$. Similarly, we have $xx^{\prime}=yy^{\prime}=zz^{\prime}=4$. Conversely, assume $xx^{\prime}=yy^{\prime}=zz^{\prime}=4$. Then, since $(xyz)^{2}=(xyz)(x^{\prime}y^{\prime}z^{\prime})=(xx^{\prime})(yy^{\prime})(zz^{\prime})=64$, we have $xyz=x^{\prime}y^{\prime}z^{\prime}=8$. Thus, we have $xyz=x^{\prime}y^{\prime}z^{\prime}=2xx^{\prime}$ and, by $x,x^{\prime}\neq 0$, it implies $y^{\prime}z^{\prime}-x=x$ and $yz-x^{\prime}=x^{\prime}$. These two equalities imply $\gamma_{1}(M)=M$. By doing a similar arguments for $\gamma_{2}(M)$ and $\gamma_{3}(M)$, we may obtain the fact that $M$ is a fixed point. ∎
When we consider $M\in\mathcal{M}_{\mathbb{Z}}^{+}$, we can write all fixed points. The matrices corresponding to the following fixed points are studied in [Lam16].
Proposition 10.9.
All fixed points of $\mathcal{M}^{+}_{\mathbb{Z}}$ for the action $\Gamma$ are
(60) $\left(\begin{matrix}2&2&2\\ 2&2&2\end{matrix}\right),\left(\begin{matrix}4&1&2\\ 1&4&2\end{matrix}\right),$
and their permutations.
Proof.
It follows from Proposition 10.9. Note that integers $z,z^{\prime}\in\mathbb{Z}_{>0}$ satisfying $zz^{\prime}=4$ are $(z,z^{\prime})=(1,4),(2,2),(4,1)$. By considering the condition $xyz=x^{\prime}y^{\prime}z^{\prime}=8$, we may show the claim. ∎
Remark 10.10.
Consider an integer skew-symmetrizable matrix $B$. Then, Proposition 10.9 implies that, for a cyclic matrix $B$, it holds that $\mu_{k}(B)=-B$ for any $k=1,2,3$ if and only if $B$ corresponds to an element in (60). However, since we may verify that, for any acyclic-matrix $B$ except for the zero matrix $O$, $\mu_{k}(B)\neq-B$ holds for some $k=1,2,3$. Thus, $\mu_{k}(B)=-B$ holds for any $k=1,2,3$ if and only if $B=O$ or $B$ corresponds to an element in (60).
The following result is a generalization of [BBH11, Cor. 1.3] for the skew-symmetrizable case.
Proposition 10.11 (cf. [BBH11, Cor. 1.3]).
All cluster-positive elements $M\in\mathcal{M}^{\mathrm{CP}}_{\mathbb{Z}}$ with $|\Gamma(M)|<\infty$ are in (60) and their permutations.
By considering the fact of [FST12b, Thm. 6.1] and [FST12a, Thm. 5.13], we may show it. However, we give an alternative proof by focusing on the partial order $\leq$.
Proof.
Assume $|\Gamma(M)|<\infty$ and set $S=\mathrm{Sk}(M)$. Then, since the map $\mathrm{Sk}:\Gamma(M)\to\Gamma(S)$ is surjective, we have $|\Gamma(S)|\leq|\Gamma(M)|<\infty$. Thus, there is a maximal element $S^{\prime}\in\Gamma(S)$. Then, since $S^{\prime}$ satisfies (M1) or (M2), $S^{\prime}\leq\gamma_{i}(S^{\prime}),\gamma_{j}(S^{\prime})$ must hold for at least two indices $i,j=1,2,3$. Without loss of generality, we may assume $i=1$ and $j=2$. Then, since $S^{\prime}$ is maximal, it also holds that $\gamma_{1}(S^{\prime}),\gamma_{2}(S^{\prime}),\gamma_{3}(S^{\prime})\leq S^{\prime}$ and, by $S^{\prime}\leq\gamma_{1}(S^{\prime}),\gamma_{2}(S^{\prime})$, we have $\gamma_{1}(S^{\prime})=\gamma_{2}(S^{\prime})=S^{\prime}$. Set $S^{\prime}=(p,q,r)$. The equalities $\gamma_{1}(S^{\prime})=\gamma_{2}(S^{\prime})=S^{\prime}$ imply $p=q$ and $r=2$. Moreover, since $\gamma_{3}(S^{\prime})=(p,p,p^{2}-2)\leq S^{\prime}$, we need $p^{2}-2p\leq 2$. Since $p\geq 2$, the inequality $p^{2}-2\leq 2$ implies $p=2$. Thus, a maximum element $M^{\prime}$ should satisfy $\mathrm{Sk}(M^{\prime})=(2,2,2)$. By Lemma 10.8, $M^{\prime}$ must be expressed as in (60). Moreover, since $\Gamma(M)=\Gamma(M^{\prime})=\{M^{\prime}\}$, $M=M^{\prime}$ holds. ∎
10.3. Surjectivity of the Markov constants
In [BBH11, Thm. 1.2 (3)], they showed that there are no integral cluster-positive elements $S$ with $40<C(S)<$ 4. Moreover, an element $S$ satisfying (M1) and $C(S)=0$ is only $S=(3,3,3)$. On the other hand, by extending integer skew-symmetrizable matrices $\mathcal{M}_{\mathbb{Z}}^{\mathrm{CP}}$ (or $\widehat{\mathcal{S}}^{\mathrm{CP}}$), we can find some elements with $40\leq C(S)<$ 4 as in Example 10.18. More strongly, we obtain the following result.
Theorem 10.12.
The maps $C:{\mathcal{M}}_{\mathbb{Z}}^{\mathrm{CP}}\to\mathbb{Z}_{\leq 4}$ and $C:\widehat{\mathcal{S}}^{\mathrm{CP}}\to\mathbb{Z}_{\leq 4}$ are surjective.
Proof.
Recall that the map $\mathrm{Sk}:\mathcal{M}\to\mathcal{S}$ satisfies $C(\mathrm{Sk}(M))=C(M)$. It implies $C({\mathcal{M}}_{\mathbb{Z}}^{\mathrm{CP}})=C(\widehat{\mathcal{S}}^{\mathrm{CP}})$. Thus, it suffices to show that $C(\widehat{\mathcal{S}}^{\mathrm{CP}})=\mathbb{Z}_{\leq 4}$. For any $n\in\mathbb{Z}_{\leq 4}$, set $S=(\sqrt{5(5-n)},2\sqrt{5-n},\sqrt{5})\in\widehat{\mathcal{S}}^{\mathrm{CP}}$. Then, we obtain $C(S)=n$. ∎
Remark 10.13.
Of course, this choice of $S$ is not unique. For example, we may take $S=(\sqrt{-n+9},\sqrt{-n+9},3)$ for $C(S)=n$.
Next, we give all elements $S\in\widehat{\mathcal{S}}^{\mathrm{CP}}$ satisfying $40\leq C(S)<$ 4 and (M1). The classification in this region is important by the following property.
Proposition 10.14 ([BBH11, Lem. 3.3]).
Let $S\in\mathcal{S}^{+}$ be not cluster-positive. Then, we have $C(S)\geq 0$. In particular, for any $S\in\mathcal{S}^{+}$, we have the following statements.
(a) If $C(S)>4$, then $S$ is not cluster-positive.
(b) If $C(S)<0$, then $S$ is cluster-positive.
For any non cluster-positive element $S\in\mathcal{S}^{+}$, there is $S^{\prime}\in\mathcal{S}$ such that two entries are positive and one entry is not positive. By considering $C(S^{\prime})$, we have $C(S)=C(S^{\prime})\geq 0$.
Thanks to Proposition 10.14, when $C(S)<0$ or $C(S)>4$, then we may determine whether $S$ is cluster-positive or not by focusing on only its Markov constant. On the other hand, when $0\leq C(S)\leq 4$, we need to consider something other than $C(S)$. In the case of $C(S)=4$, we give a condition in Lemma 8.2.
Later, we give a method to obtain all orbits satisfying $C(S)=C$ with $C<4$. Recall that, for any cluster-positive elemenet $S\in\widehat{\mathcal{S}}^{\mathrm{CP}}$ corresponding to an integer skew-symmetrizable matrix, its orbit $\Gamma(S)$ may be represented by the (M1) element (Theorem 10.7).
Lemma 10.15.
Let $S=(p,q,r)\in{\mathcal{S}}^{+}$ satisfy (M1) and $p\geq q\geq r$ and let $C\in\mathbb{Z}_{<4}$.
(a) There exists a real number $R>2$ satisfying $C(R,R,R)=C$.
(b) If $C(S)\geq C$, then we have $r\leq R$, where $R$ is the real number satisfying $C(R,R,R)=C$.
Proof.
(a) Since $\frac{d}{dx}C(x,x,x)=3x(2-x)<0$ for $x>2$, $C(x,x,x)$ is monotonically decreasing for $x\geq 2$. Moreover, since $C(2,2,2)=4$ and $\lim_{x\to\infty}C(x,x,x)=-\infty$, there is $R>2$ satisfying $C(R,R,R)=C$ uniquely.
(b) Since $S$ satisfies (M1), we have $r\geq 2$. By Lemma 8.2 (1), we have
(61) $C(S)$ $\leq C(r,r,r).$
Since $C=C(R,R,R)\leq C(S)\leq C(r,r,r)$, we have $C(R,R,R)\leq C(r,r,r)$. Moreover, since $C(x,x,x)$ is monotonically decreasing, we have $r\leq R$. ∎
Lemma 10.16.
Let $S=(p,q,r)\in\mathcal{S}^{+}$ satisfy (M1) and $p\geq q\geq r$. Let $C\in\mathbb{Z}_{<4}$. If $C=C(S)$, then we have $2<r$ anf $p\leq r\sqrt{\frac{r^{2}-C}{r^{2}-4}}$.
Proof.
Since $C(S)=C<4$, we have $r>2$. (If $r=2$, we have $C(S)=(p-q)^{2}+4\geq 4$.) By solving the equation $C(S)=p^{2}+q^{2}+r^{2}-pqr=C$ for $q$, we have $q=\frac{1}{2}(pr\pm\sqrt{p^{2}(r^{2}-4)-4(r^{2}-C)})$. Note that $\sqrt{p^{2}(r^{2}-4)-4(r^{2}-C)}$ is a real number since $q\in\mathbb{R}$. If we choose $q=\frac{1}{2}(pr+\sqrt{p^{2}(r^{2}-4)-4(r^{2}-C)})$, then we have $q\geq\frac{1}{2}pr>p$ by $r>2$. Thus, since $q\leq p$, we should choose $q=\frac{1}{2}(pr-\sqrt{p^{2}(r^{2}-4)-4(r^{2}-C)})$. Moreover, since $S$ satisfies (M1), we have $q\geq\frac{2}{r}p$ by Proposition 7.3 (1). Thus, it holds that
(62) $\frac{1}{2}(pr-\sqrt{p^{2}(r^{2}-4)-4(r^{2}-C)})\geq\frac{2}{r}p.$
It implies that $(\frac{r}{2}-\frac{2}{r})p\geq\frac{1}{2}\sqrt{p^{2}(r^{2}-4)-4(r^{2}-C)}$. Since both sides are positive, it implies that
(63) $\left(\frac{r}{2}-\frac{2}{r}\right)^{2}p^{2}\geq\frac{1}{4}\{p^{2}(r^{2}-4)-4(r^{2}-C)\},$
and we have
(64) $r^{2}-C\geq\frac{r^{2}-4}{r^{2}}p^{2}.$
Since $r^{2}-4,r^{2}>0$, we have $p^{2}\leq r^{2}\frac{r^{2}-C}{r^{2}-4}$. Note that $r^{2}-C\geq 0$ since $r^{2}\geq 4\geq C=C(S)$ by Lemma 8.1. Thus, this inequality implies that $p\leq r\sqrt{\frac{r^{2}-C}{r^{2}-4}}$. ∎
Proposition 10.17.
Let $S=(p,q,r)\in\widehat{\mathcal{S}}^{\mathrm{CP}}$ satisfy (M1) and $p\geq q\geq r$ and $C\in\mathbb{Z}_{<4}$. Set $R>2$ such that $C(R,R,R)=C$. If $C(S)=C$, then we have
(65) $2<r\leq R,\quad q\leq p\leq r\sqrt{\frac{r^{2}-C}{r^{2}-4}}.$
In particular, for each $C\in\mathbb{Z}_{<4}$, the number of possible $S\in\widehat{\mathcal{S}}^{\mathrm{CP}}$ satisfying (M1) and $C(S)=C$ is finite.
Example 10.18.
Consider an element $S=(p,q,r)\in\widehat{\mathcal{S}}^{\mathrm{CP}}$ satisfying (M1) and $40\leq C(S)<$ 4. The real number $R>2$ satisfying $C(R,R,R)=3R^{2}-R^{3}=0$ is $R=3$. Thus, by Lemma 10.15, under the assumption $p\geq q\geq r$, we need $2<r\leq 3$. Since $r^{2}\in\mathbb{Z}$, we have $\sqrt{5}\leq r\leq 3$. Moreover, by setting $C=C(S)$, we have $p\leq r\sqrt{\frac{r^{2}-C}{r^{2}-4}}\leq 3\times\sqrt{\frac{3^{2}-0}{\sqrt{5}^{2}-4}}=9$ by Lemma 10.16. In particular, by considering $p^{2},q^{2},r^{2}\in\mathbb{Z}$, the number of possible $S$ are at most $5\times 81^{2}$. (By considering other conditions such as $pqr\in\mathbb{Z}$ and $r\leq q\leq p$, the number of elements which we need to check is much smaller than $5\times 81^{2}$.) We may write all cluster-positive orbits satisfying $40\leq C(S)<$ 4 as follows:
Proposition 10.19.
For any element $S\in\widehat{\mathcal{S}}$ corresponding to an integer skew-symmetrizable matrix,
(a) $S$ is cluster-positive with $C(S)=0$ if and only if $S$ belongs to any of the following sets or their permutations.
(66) $\Gamma(5,2\sqrt{5},\sqrt{5}),\Gamma(3\sqrt{2},2\sqrt{3},\sqrt{6}),\Gamma(4,2\sqrt{2},2\sqrt{2}),\Gamma(3,3,3).$
(b) $S$ is cluster-positive with $C(S)=1$ if and only if $S$ belongs to any of the following sets or their permutations.
(67) $\Gamma(2\sqrt{5},4,\sqrt{5}),\Gamma(\sqrt{15},\sqrt{10},\sqrt{6}),\Gamma(\sqrt{14},2\sqrt{2},\sqrt{7}),\Gamma(3,2\sqrt{2},2\sqrt{2}).$
(c) $S$ is cluster-positive with $C(S)=2$ if and only if $S$ belongs to any of the following sets or their permutations.
(68) $\Gamma(\sqrt{15},2\sqrt{3},\sqrt{5}),\Gamma(2\sqrt{3},2\sqrt{2},\sqrt{6}),\Gamma(3,\sqrt{7},\sqrt{7}).$
(d) $S$ is cluster-positive with $C(S)=3$ if and only if $S$ belongs to any of the following sets or their permutations.
(69) $\Gamma(\sqrt{10},2\sqrt{2},\sqrt{5}),\Gamma(3,\sqrt{6},\sqrt{6}).$
(e) $S$ is cluster-positive with $C(S)=4$ if and only if $S$ belongs to $\Gamma(p,p,2)$ for some $p\geq 2$ with $p^{2}\in\mathbb{Z}$ or their permutations.

### v2

@@
 In the usual cluster algebra theory, we often suppose that all entries of $B$ are integral, since entries of $B$ appear in the exponent of rational fractions.
-However, since we do not focus on cluster variables, we consider real skew-symmetrizable matrices.
+However, since we do not focus on cluster variables, we may consider real skew-symmetrizable matrices.
 When we focus on only matrix mutations $\mu_{k}(B)$, we can naturally extend the definition, and their properties are helpful for the study of cluster algebras.
@@
 Our purpus is to identify whether $B$ is cluster-cyclic or not.
+1.2. Previous works
 In the skew-symmetric case, this problem was solved in [BBH11, FT19].
@@
 For the skew-symmetrizable case, the Markov constants $C(B)=xx^{\prime}+yy^{\prime}+zz^{\prime}-|xyz|$, which are a generalization of the one for skew-symmetric matrices, also play important roles in this problem.
+For the skew-symmetrizable case, this problem was solved by Seven [Sev12] by using a quasi-Cartan companion.
 Other motivation of this study is in the classification of the global structures of $G$-fans [Nak24].
+Remark 1.1.
+After the author submitted this paper, I was informed by Ahmet Seven of the related works in [Sev12, Sev13].
+By using results in [Sev12], Theorem 1.2 can be shown.
+A related result to Theorem 9.7 appeared in [Sev13].
+However, to obtain other results, we necessitate a different proof that draws upon certain facts from the proof of Theorem 1.2.
+So, we give an alternative proof of Theorem 1.2 and Theorem 9.7 in this paper.
 1.3. Method
@@
 One of the main reasons is that the surjectivity of Markov constants.
-(See Theorem 10.12, Example 10.18.) In [BBH11], it was shown that there are no integer skew-symmetric and cluster-cyclic matrices with $40<C(B)<$ 4.
-Thanks to this property, some arguments became clear.
-However, by considering skew-symmetrizable matrices, there are both cluster-cyclic and cluster-acyclic matrices $B$ with $40<C(B)<$ 4.
+(See Theorem 9.12, Example 9.18.) In [BBH11], it was shown that there are no integer skew-symmetric and cluster-cyclic matrices with $0<C(B)<4$.
+This property simplifies certain arguments compared to the skew-symmetrizable case.
+However, due to this difference, we require some modifications to its proof.
 1.4. Main results
-First, we mention the identification of cluster-cyclicity. It is a natural generalization of [BBH11] for the skew-symmetrizable and real case.
-Theorem 1.1 (cf. [BBH11, Thm. 1.2 (1),(2)]).
-Let $B$ be a skew-symmetrizable and cyclic matrix of the form (1). Then, the following two conditions are equivalent.
+First, we mention the identification of cluster-cyclicity.
+It is a natural generalization of [BBH11] for the skew-symmetrizable and real case.
+The following result has already obtained in [Sev12] essentially by using a quasi-Cartan companion.
+Theorem 1.2 (cf. [BBH11, Thm. 1.2 (1),(2)], [Sev12, Thm. 2.6]).
+Let $B$ be a skew-symmetrizable and cyclic matrix of the form (1). Define its Markov constant
+(2) $C(B)=xx^{\prime}+yy^{\prime}+zz^{\prime}-|xyz|.$
+Then, the following two conditions are equivalent.
   • 1.
@@
 By considering $\gamma_{k}$, if $B$ is positive and cluster-cyclic, then every element in this orbit $\Gamma(B)$ is also so.
-(See Proposition 4.1.)
-Next result is about a fundamental domain for this action $\Gamma$ of integer skew-symmetrizable matrices.
+Next result is to give a fundamental domain for this action $\Gamma$ of integer skew-symmetrizable matrices.
 The following partial order $\leq$ on skew-symmetrizable and positive-cyclic matrices is useful in this study.
@@
 A set $\tilde{F}$ defined by
-(3) $\left\{B=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\\ \end{matrix}\right)\mid\begin{aligned} &\textup{$B$ is skew-symmetrizable},\\ &xyz\geq 2xx^{\prime},2yy^{\prime},2zz^{\prime},\\ &x,y,z,x^{\prime},y^{\prime},z^{\prime}\in\mathbb{Z}_{>0}\end{aligned}\right\}$
+(4) $\left\{\left.B=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\\ \end{matrix}\right)\ \right|\ \begin{aligned} &\textup{$B$ is skew-symmetrizable},\\ &xyz\geq 2xx^{\prime},2yy^{\prime},2zz^{\prime},\\ &x,y,z,x^{\prime},y^{\prime},z^{\prime}\in\mathbb{Z}_{>0}\end{aligned}\right\}$
 is a fundamental domain for $\Gamma$ on integer skew-symmetrizable, positive and cluster-cyclic matrices.
 Moreover, for any integer skew-symmetrizable, positive, and cluster-cyclic matrix $B$, the element of $\Gamma(B)\cap\tilde{F}$ is the smallest element in $\Gamma(B)$.
-Nemely, for each integer skew-symmetrizable, positive, and cluster-cyclic matrix $B$, there is the minimum element of $\Gamma(B)$ and it is an element in $\tilde{F}$.
+Namely, for each integer skew-symmetrizable, positive, and cluster-cyclic matrix $B$, there is the minimum element of $\Gamma(B)$, and $\tilde{F}$ is the set of all such minimum elements.
 Unfortunately, this result cannnot generalize for the real case.
@@
 These two results are natural generalizations of the skew-symmetric case.
-On the other hand, the following result is the most important difference between the skew-symmetrizable case and the skew-symmetric case.
+On the other hand, the following result is the biggest difference between the skew-symmetrizable case and the skew-symmetric case.
 Theorem 1.4 (Theorem 9.12).
-The Markov constant $C(B)$ can take any integer less than or equal to $4$ by choosing an integer skew-symmetrizable and cluster-cyclic matrix $B$.
-Because of this property (in particular, the fact that there is a cluster-cyclic matrix $B$ with $40<C(B)<$ 4), we need to reconstruct a proof in [BBH11] even though we focus on the integer skew-symmetrizable matrices (see Example 10.18).
+The Markov constant $C(B)$ can take any integer less than or equal to 444 by choosing an integer skew-symmetrizable and cluster-cyclic matrix $B$.
+[…]
 1.5. Structure of the paper
@@
 In section 3, we define the Markov constant. This is a generalization for the skew-symmetrizable case of the one which is introduced in [BBH11].
-In section 4, we introduce more simple notations, which is similar to [BBH11, FT19]. In this notations, Theorem 1.1 is decomposed to Lemma 4.7, Lemma 4.8, and Lemma 4.9.
+In section 4, we introduce more simple notations, which is similar to [BBH11, FT19]. In this notations, we may decompose Theorem 1.2 into Lemma 4.7, Lemma 4.8, and Lemma 4.9.
 In section 5, we show Lemma 4.7.
-In section 6, we calculate some examples and prove Lemma 4.8. Most contents and ideas were in [FT19].
+In section 6, we show that some mutation can be expressed by the Chebyshev polynomials of the second kind. Based on this, we may prove Lemma 4.8. These ideas have appeared in [FT19].
 In section 7 and 8, we show Lemma 4.9. Most ideas are similar to the one in [BBH11]. However some modifications are necessarily to generalize it for the real and skew-symmetrizable case.
 In section 9 and 10, we derive some new results.
-In section 9, we consider the case which does not occur when we consider integer skew-symmetric matrces.
-The related topic is in [FT19].
 In section 9, we deal with the integer skew-symmetrizable case.
 In particular, we focus on the difference between the result of the skew-symmetric case in [BBH11] and the one of the skew-symmetrizable case.
+In section 10, we consider the case which does not occur when we consider integer skew-symmetric matrces.
 Acknowledgement.
 This problem was suggested by Tomoki Nakanishi, and I thank him for various important advices. I also thank Zhichao Chen for careful reading and advices.
+I thank Ahmet Seven for noticing the important related works.
 This work was financially supported by JST SPRING, Grant Number JPMJSP2125.
@@
 2.1. Skew-symmetrizable matrix and mutation
-In this paper, we will focus on skew-symmetrizable matrices of rank $3$ with real entries.
-Let $\mathrm{M}_{3}(\mathbb{R})$ be the set of all $3\times 3$ real matrices.
+In this paper, we will focus on skew-symmetrizable matrices of rank 333 with real entries.
+Let $\mathrm{M}_{3}(\mathbb{R})$ be the set of all $\times 3$ 3 real matrices.
 A real matrix $B=(b_{ij})\in\mathrm{M}_{3}(\mathbb{R})$ is called skew-symmetrizable if there exists a diagonal matrix $D=\mathrm{diag}(d_{1},d_{2},d_{3})$ with positive real numbers $d_{1},d_{2},d_{3}>0$ such that $DB$ is skew-symmetric.
@@
 Let $B=(b_{ij})\in\mathrm{M}_{3}(\mathbb{R})$ be a skew-symmetrizable matrix. Then, we have the following statements.
-(a) $B$ is sign-skew-symmetric, namely, $b_{ij}$ and $-b_{ji}$ have the same sign for any $i,j$. (In particular, every diagonal entry is 00.)
+(a) $B$ is sign-skew-symmetric, namely, $b_{ij}$ and $-b_{ji}$ have the same sign for any $i,j$. (In particular, every diagonal entry is 0.)
 (b) We have
-(4) $b_{12}b_{23}b_{31}=-b_{21}b_{32}b_{13}.$
+(5) $b_{12}b_{23}b_{31}=-b_{32}b_{21}b_{13}.$
 Conventionally, the sign $\varepsilon_{x}$ of a real number $x\in\mathbb{R}$ is defined as follows: $\varepsilon_{x}=0$ if $x=0$ and, otherwise, $\varepsilon_{x}=\pm 1$ with $\varepsilon_{x}x>0$.
@@
 For any skew-symmetrizable matrix $B$, we may define the corresponding quiver $Q(B)$ wighted by $\mathbb{R}_{\geq 0}^{2}$ as follows:
-  • The vertices of $Q(B)$ are indexed by $1$, $2$, and $3$.
+  • The vertices of $Q(B)$ are indexed by 111, 222, and 333.
   • For each vertex pair $i\neq j$, if $b_{ij}>0$ or, equivalently, $b_{ji}<0$, there is an edge $i\overset{(b_{ij},-b_{ji})}{\longrightarrow}j$.
@@
 The corresponding quivers are described as in Figure 2 and Figure 2.
-111122223333$(y,y^{\prime})$$(x,x^{\prime})$$(z,z^{\prime})$
+[…]
 Figure 1. Positive
-111122223333$(-y^{\prime},-y)$$(-x^{\prime},-x)$$(-z^{\prime},-z)$
+[…]
 Figure 2. Negative
@@
 Definition 3.1.
-For any skew-symmetrizable matrix $B\in\mathrm{M}_{3}(\mathbb{R})$ of the form (13), we define the Markov constant $C(B)$ of $B$ as follows:
+For any cyclic matrix $B\in\mathrm{M}_{3}(\mathbb{R})$ of the form (14), define the Markov constant
 (15) $C(B)=xx^{\prime}+yy^{\prime}+zz^{\prime}-|xyz|.$
-Similarly, we define $C^{+}(B)$ and $C^{-}(B)$ as
-(15) $C^{\pm}(B)=xx^{\prime}+yy^{\prime}+zz^{\prime}-(\pm xyz).$
-These three numbers have the following relations.
+The following properties are fundamental.
 Proposition 3.2.
 Let $B$ be a skew-symmetrizable matrix.
-(a) We have $C(B)=C(-B)$ and $C^{+}(B)=C^{-}(-B)$.
-(b) If $B$ is positive-cyclic (resp. negative-cyclic), then we have $C(B)=C^{+}(B)$ (resp. $C(B)=C^{-}(B)$).
-We may obtain these properties by considering a sign of $xyz$.
-For any skew-symmetrizable matrix $B$, since $xyz=x^{\prime}y^{\prime}z^{\prime}=pqr$ as shown in (10), the Markov constant $C(B)$ may also be expressed as follows:
-(16) $C(B)$ $=xx^{\prime}+yy^{\prime}+zz^{\prime}-|xyz|$
-$=xx^{\prime}+yy^{\prime}+zz^{\prime}-|x^{\prime}y^{\prime}z^{\prime}|$
-$=p^{2}+q^{2}+r^{2}-|pqr|=C(\mathrm{Sk}(B)).$
-The following proposition is fundamental for later.
-Proposition 3.3.
-Let $B$ be a skew-symmetrizable matrix. Then, the following two statements hold.
-(a) We have $C^{\pm}(B)=C^{\pm}(\mathrm{Sk}(B))$ and $C(B)=C(\mathrm{Sk}(B))$.
-(b) If $B$ is positive-cyclic (resp.
-negative-cyclic), we have $C^{-}(\mu_{k}(B))=C^{+}(B)$ (resp.
-$C^{+}(\mu_{k}(B))=C^{-}(B)$) for any $k=1,2,3$.
-In particular, we have $C^{+}(-\mu_{k}(B))=C^{+}(B)$ if $B$ is positive-cyclic.
+(a) We have $C(\mathrm{Sk}(B))=C(B)$.
+(b) For any $k=1,2,3$, we have $C(\mu_{k}(B))=C(B)$.
 Proof.
-Let $B$ be expressed as in (13).
-The statement (a) has already been shown in (16).
-We show (b) in the case of $k=1$.
-Let $\varepsilon$ be $+$ and -- if $B$ is positive-cyclic and negative-cyclic, respectively.
-Since $C^{\varepsilon}(B)=C^{\varepsilon}(\mathrm{Sk}(B))$ and $C^{-\varepsilon}(\mu_{1}(B))=C^{-\varepsilon}(\mathrm{Sk}(\mu_{1}(B)))=C^{-\varepsilon}(\mu_{1}(\mathrm{Sk}(B)))$ by (a) and Proposition 2.6 (b), it suffices to show that $C^{\varepsilon}(\mathrm{Sk}(B))=C^{-\varepsilon}(\mu_{1}(\mathrm{Sk}(B)))$.
-We have
-(17) $\mu_{1}(\mathrm{Sk}(B))=\left(\begin{matrix}0&r&-q\\ -r&0&-p+\varepsilon qr\\ q&p-\varepsilon qr&0\end{matrix}\right).$
-Thus, we obtain
-(18) $C^{-\varepsilon}(\mu_{1}(\mathrm{Sk}(B)))$ $=(p-\varepsilon qr)^{2}+q^{2}+r^{2}-(-\varepsilon)(p-\varepsilon qr)(-q)(-r)$
-$=p^{2}+q^{2}+r^{2}-\varepsilon pqr=C^{\varepsilon}(\mathrm{Sk}(B)).$
-4. Reduction of Theorem 1.1 to the skew-symmetric case
+The statement (a) follows from (11). We show (b). Set
+(16) $B=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\end{matrix}\right),\ \mu_{k}(B)=\left(\begin{matrix}0&w^{\prime}&-v\\ -w&0&u^{\prime}\\ v^{\prime}&-u&0\end{matrix}\right).$
+Suppose that $B$ is positive-cyclic. Regardless $B$ and $\mu_{k}(B)$ is cyclic or not, we have $\varepsilon_{\mu_{k}(B)}|uvw|=uvw$. Thus, we have
+(17) $C(\mu_{k}(B))=uu^{\prime}+vv^{\prime}+ww^{\prime}-uvw.$
+We can check $C(\mu_{k}(B))=C(B)$ by a direct calculation. For example, if $k=1$, we have
+(18) $C(\mu_{1}(B))$ $=(y^{\prime}z^{\prime}-x)(yz-x^{\prime})+yy^{\prime}+zz^{\prime}-(y^{\prime}z^{\prime}-x)yz$
+$=xx^{\prime}+yy^{\prime}+zz^{\prime}-x^{\prime}y^{\prime}z^{\prime}=C(B).$
+Suppose that $B$ is acyclic.
+For the simplicity, we assume $x,x^{\prime}\leq 0$ and $y,y^{\prime},z,z^{\prime}>0$.
+Then, regardless of the cyclicity of $\mu_{k}(B)$, we have $\varepsilon_{\mu_{k}(B)}|uvw|=-uvw$ holds.
+By a similar calculation of (18), we may show $C(\mu_{k}(B))=C(B)$.
+∎
+4. Notations
 In this section, we give more simple notations as in [BBH11], and we translate Theorem 1.2 into these notations.
-Let $B\in\mathrm{M}_{3}(\mathbb{R})$ be a skew-symmetrizable matrix.
-Let $t=(t_{1},t_{2},\dots,t_{l})$ be a finite sequence of indices $1,2,3$ such that $t_{i}\neq t_{i+1}$ for any $i=1,2,\dots,l-1$.
-The length $|t|$ is defined by the number of indices; in this case, the length of $t$ is $l$.
-We write $\mu_{t}(B)=\mu_{t_{l}}\cdots\mu_{t_{2}}\mu_{t_{1}}(B)$.
-A collection $\Sigma(B)=\{\mu_{t}(B)\}_{t}$ indexed by all of such $t$ is called the matrix-pattern of $B$.
-In convention, we allow $t=\emptyset$ and let $\mu_{\emptyset}=\mathrm{id}$.
-We often fix one element of $\Sigma(B)$ and we call it an initial matrix of $\Sigma(B)$.
-Now, we may see $B$ as an initial marix of $\Sigma(B)$.
-Then the distance (from $B$) to $\mu_{t}(B)\in\Sigma(B)$ is defined by the length of $t$.
-Let
-(19) $B=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\end{matrix}\right)$
-be a skew-symmetrizable and cyclic matrix.
-Recall that $B$ is said to be positive-cyclic (resp.
-negative-cyclic) if all of $x,y,z,x^{\prime},y^{\prime},z^{\prime}$ are positive (resp.
-negative).
-Later, we often consider positive-cyclic matrices.
-If we consider a negative-cyclic matrix $B$, apply the results for $-B$.
-Our interest is whether a given positive-cyclic matrix $B\in\mathrm{M}_{3}(\mathbb{Z})$ is cluster-cyclic or not.
-In this case, it is enough to consider the following maps $\gamma_{k}$ instead of $\mu_{k}$.
-(20) $\gamma_{1}(B)$ $=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-(yz-x^{\prime})\\ -y^{\prime}&y^{\prime}z^{\prime}-x&0\end{matrix}\right),$
-$\gamma_{2}(B)$ $=\left(\begin{matrix}0&-z^{\prime}&z^{\prime}x^{\prime}-y\\ z&0&-x^{\prime}\\ -(zx-y^{\prime})&x&0\end{matrix}\right),$
-$\gamma_{3}(B)$ $=\left(\begin{matrix}0&-(xy-z^{\prime})&y\\ x^{\prime}y^{\prime}-z&0&-x^{\prime}\\ -y^{\prime}&x&0\end{matrix}\right),$
-where the domain of these functions consists of all skew-symmetrizable matrices (not necessarily cyclic).
-We may verify that $\gamma_{k}\gamma_{k}(B)=B$.
-Define $\Gamma(B)=\{\gamma_{t}(B)\}_{t}$ in the same way as $\Sigma(B)$, but we replace $\mu_{k}$ with $\gamma_{k}$.
-In this case, we have the following proposition.
-Proposition 4.1.
-Let $B$ be a skew-symmetrizable and positive-cyclic matrix.
-(a) We have $\gamma_{k}(B)=-\mu_{k}(B)$ for any $k=1,2,3$.
-(b) The following two conditions are equivalent.
-  • (1)
-    $B$ is cluster-cyclic.
-  • (2)
-    Every matrix of $\Gamma(B)$ is positive-cyclic.
-Proof.
-We may show (a) by the direct calculation.
-By considering (a), we may show (b) as follows: Note that $\mu_{k}(-B)=-\mu_{k}(B)$.
-For a sequence of indices $t=(t_{1},\dots,t_{l})$ and $t^{\prime}=(t_{1},\dots,t_{l},t_{l+1})$, if $\gamma_{t}(B)=(-1)^{|t|}\mu_{t}(B)$ is cluster-cyclic, we have $\gamma_{t^{\prime}}(B)=(-1)^{|t^{\prime}|}\mu_{t^{\prime}}(B)$.
-This relation is true when $\mu_{t}(B)$ is cyclic.
-(we can disregard the condition of whether $\mu_{t^{\prime}}(B)$ is cyclic or not.) Thus, if there is an acyclic-matrix $\mu_{t^{\prime}}(B)$ and $|t^{\prime}|$ is the smallest among such $t^{\prime}$, then it holds that $\gamma_{t^{\prime}}(B)=(-1)^{|t^{\prime}|}\mu_{t^{\prime}}(B)$ and it is acyclic.
-∎
-Later, we write a skew-symmetrizable matrix by a following tuplet:
+Our purpose is to identify whether a given skew-symmetrizable matrix $B$ is cluster-cyclic or not.
+In this case, $\gamma_{k}=-\mu_{k}$ is more convenient than the usual mutation $\mu_{k}$.
+Note that $-\mu_{k}(B)=\mu_{k}(-B)$ for any skew-symmetrizable matrix $B$.
+Thus, $B$ is cluster-cyclic if and only if every $\gamma_{i_{s}}\cdots\gamma_{i_{1}}(B)=(-1)^{s}\mu_{i_{s}}\cdots\mu_{i_{1}}(B)$ is cyclic.
+Moreover, for any skew-symmetrizable matrix
+(19) $B=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\end{matrix}\right),$
+we will write it as the following tuplet:
 (20) $M=\left(\begin{matrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{matrix}\right).$
@@
 and the sign of two entries on the same row is the same.
-The maps $\gamma_{k}$ may be expressed as
+Moreover, $B$ is cyclic if and only if all entries of $M$ have the same sign. If $B$ is cyclic, the maps $\gamma_{k}$ may be expressed as
 (22) $\gamma_{1}(M)$ $=\left(\begin{matrix}y^{\prime}z^{\prime}-x&y&z\\ yz-x^{\prime}&y^{\prime}&z^{\prime}\end{matrix}\right),$
@@
 (23) $\mathrm{Sk}(B)=\left(\begin{matrix}0&-r&q\\ r&0&-p\\ -q&p&0\end{matrix}\right)$
-is cluster-cyclic or not. We may see this matrix as a triplet of real numbers $(p,q,r)\in\mathbb{R}^{3}$ and the maps $\gamma_{k}$ become
+is cluster-cyclic or not. We may see this matrix as a triplet of real numbers $(p,q,r)\in\mathbb{R}^{3}$. If $\mathrm{Sk}(B)$ is cyclic, the maps $\gamma_{k}$ may be expressed as
 (24) $\gamma_{1}(p,q,r)$ $=(qr-p,q,r),$
@@
 $\gamma_{3}(p,q,r)$ $=(p,q,pq-r).$
+Our interest is for cyclic matrices, and we do not treat acyclic matrixes. So, we see (22) and (24) as the definition of $\gamma_{k}$.
 We summerize the definitions.
@@
 […]
-+=$\mathcal{M}^{+}=\left\{\left(\begin{matrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{matrix}\right)\in\mathcal{M}\mid x,y,z,x^{\prime},y^{\prime},z^{\prime}>0\right\},$
-$\mathcal{S}=\{(p,q,r)\mid p,q,r\in\mathbb{R}\}=\mathbb{R}^{3},$
-$\mathcal{S}^{+}=\{(p,q,r)\in\mathcal{S}\mid p,q,r>0\}.$
+$\mathcal{M}^{+}$ =$=\left\{\left.\left(\begin{matrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{matrix}\right)\in\mathcal{M}\ \right|\ x,y,z,x^{\prime},y^{\prime},z^{\prime}>0\right\},$
+$\mathcal{S}$ =$=\{(p,q,r)\mid p,q,r\in\mathbb{R}\}=\mathbb{R}^{3},$
+$\mathcal{S}^{+}$ =$=\{(p,q,r)\in\mathcal{S}\mid p,q,r>0\}.$
 Every element of $\mathcal{M}^{+}$ and $\mathcal{S}^{+}$ is called positive.
 We define maps $\gamma_{k}:\mathcal{M}\to\mathcal{M}$ and $\gamma_{k}:\mathcal{S}\to\mathcal{S}$ as (22) and (24), respectively.
-The maps $\gamma_{k}$ satisfy $\gamma_{k}^{2}=\mathrm{id}$, in particular, $\gamma_{k}$ is a bijection on $\mathcal{S}$.
 Definition 4.2.
-For any $M\in\mathcal{M}$, let $\Gamma(M)$ be the collection of $\gamma_{t}(M)$ indexed by all of finite sequences $t=(t_{1},\dots,t_{k})$ of indices $1,2,3$ with $t_{i}\neq t_{i+1}$, where $\gamma_{t}=\gamma_{t_{k}}\cdots\gamma_{t_{1}}$.
+For any finite sequence $t=(t_{1},\dots,t_{k})$ of indices 1,2,31,2,31 , 2 , 3, set $\gamma_{t}=\gamma_{t_{k}}\cdots\gamma_{t_{1}}$.
+For any $M\in\mathcal{M}$, let $\Gamma(M)$ be the collection of $\gamma_{t}(M)$ indexed by all of finite sequences $t=(t_{1},\dots,t_{k})$ of indices 1,2,31,2,31 , 2 , 3 with $t_{i}\neq t_{i+1}$, where $\gamma_{t}=\gamma_{t_{k}}\cdots\gamma_{t_{1}}$.
 For any positive element $M\in\mathcal{M}^{+}$, if every element in $\Gamma(M)$ is positive, $M$ is said to be cluster-positive.
 Similarly, we define the cluster-positivity of $S\in\mathcal{S}^{+}$.
-As we said, every positive element $M\in\mathcal{M}^{+}$ and $S\in\mathcal{S}^{+}$ corresponds to a positive-cyclic matrices. Here, we define a map $\mathrm{Sk}:\mathcal{M}\to\mathcal{S}$ as
+Here, we define a map $\mathrm{Sk}:\mathcal{M}\to\mathcal{S}$ as
 (26) $\mathrm{Sk}\left(\begin{matrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{matrix}\right)=(\varepsilon_{x}\sqrt{xx^{\prime}},\varepsilon_{y}\sqrt{yy^{\prime}},\varepsilon_{z}\sqrt{zz^{\prime}}),$
 where $\varepsilon_{a}\in\{0,\pm 1\}$ is the sign of $a\in\mathbb{R}$.
-Obviously, if $M\in\mathcal{M}$ corresponds to a positive-cyclic matrix $B\in\mathrm{M}_{3}(\mathbb{R})$, then $\mathrm{Sk}(M)$ corresponds to its double-sided skew-symmetrized form $\mathrm{Sk}(B)$.
+If $M\in\mathcal{M}$ corresponds to a positive-cyclic matrix $B\in\mathrm{M}_{3}(\mathbb{R})$, then $\mathrm{Sk}(M)$ corresponds to its double-sided skew-symmetrized form $\mathrm{Sk}(B)$.
 Moreover, $M$ is cluster-positive if and only if $S=\mathrm{Sk}(M)$ is cluster-positive.
@@
 By definition, we have $C(\mathrm{Sk}(M))=C(M)$. Note that, for any positive-cyclic and skew-symmetrizable matrix $B$, it corresponds to an element $M\in\mathcal{M}^{+}$. Then, it holds that
-(28) $C(M)=C^{+}(B)=C(B).$
-Moreover, we have
+(27) $C(M)=C(B).$
+Moreover, we may check
 (28) $C(\gamma_{k}(M))=C(M).$
-Proof is similar to the one of Proposition 3.3 (b), in particular, (18). Thanks to this result, the Markov constant does not change under the transformations of $\gamma_{k}$.
 We summarize the results in this section.
@@
 be skew-symmetrizable and positive-cyclic. Set $p=\sqrt{xx^{\prime}}$, $q=\sqrt{yy^{\prime}}$, $r=\sqrt{zz^{\prime}}$, and $S=(p,q,r)\in\mathcal{S}^{+}$.
-(Proposition 4.1) $B$ is cluster-cyclic if and only if $S$ is cluster-cyclic.
+(Proposition 2.6) $B$ is cluster-cyclic if and only if $S$ is cluster-positive.
 (27) We have $C(B)=C(S)$.
@@
 Proposition 4.5 ([FT19, Lem. 3.3, Thm. 4.5, Cor. 4.8]).
-Let $S=(p,q,r)\in\mathcal{S}^{+}$. Then, $S$ is cluster-cyclic if and only if both $p,q,r\geq 2$ and $C(S)\leq 4$ hold.
+Let $S=(p,q,r)\in\mathcal{S}^{+}$. Then, $S$ is cluster-positive if and only if both $p,q,r\geq 2$ and $C(S)\leq 4$ hold.
+Remark 4.6.
+This result has essentially generalized for skew-symmetrizable case by [Sev12]. In this paper, the result is expressed by the quasi-Cartan companion instead of the Markov constant.
 Now, we show Theorem 1.2 under the assumption of Proposition 4.5.
 Proof of Theorem 1.2.
-Let $B$ be a skew-symmetrizable and cyclic matrix.
-If $B$ is positive-cyclic, the claim holds by using the correspondence of $\mathrm{Sk}(B)$ and $S$.
-When $B$ is negative-cyclic, then $-B$ is positive-cyclic.
-Thus, by applying the result of Proposition 4.6 for $-B$, we can show Theorem 1.1.
-∎
+[…]
 Proposition 4.5 was shown in [FT19] by using geometrical aspects of mutations.
@@
 5. Proof of Lemma 4.7
-Here, we show Lemma 4.7. The essential part is as follows.
+Here, we show Lemma 4.7. The essential part is the following.
 Lemma 5.1.
@@
 By definition of $\gamma_{k}$, all entries other than $k$th one do not change.
-In particular, they are larger or equal to $2$.
+In particular, they are larger or equal to 222.
 By Proposition 4.4 (c), $C(S^{\prime})=C(S)\leq 4$ holds.
-Thus, the $k$th entry of $S^{\prime}$ is also larger or equal to $2$ by Lemma 5.1.
-By repeating this argument, every entry of elements in $\Gamma(S)$ is larger or equal to $2$ and, in particular, positive.
+Thus, the $k$th entry of $S^{\prime}$ is also larger or equal to 222 by Lemma 5.1.
+By repeating this argument, every entry of elements in $\Gamma(S)$ is larger or equal to 222 and, in particular, positive.
 Thus, the claim holds.
 ∎
-6. $1,2$-orbits and proof of Lemma 4.8
-This contents was introduced in [FT19, Lem. 3.3].
+6. 1,21,21 , 2-orbits and proof of Lemma 4.8
+The contents in this section have already appeared in [FT19, Lem. 3.3].
 Definition 6.1.
@@
 Definition 6.2.
-Let $S\in\mathcal{S}^{+}$. We define a collection $\Gamma_{1,2}(S)=\{\gamma_{t}(S)\}_{t}$ indexed by a sequence $t=(t_{1},\dots,t_{l})$ of $1,2$ such that $t_{i}\neq t_{i+1}$ for any $i=1,\dots,l-1$.
+Let $S\in\mathcal{S}^{+}$.
+We define a collection $\Gamma_{1,2}(S)=\{\gamma_{t}(S)\}_{t}$ indexed by a sequence $t=(t_{1},\dots,t_{l})$ of 1,21,21 , 2 such that $t_{i}\neq t_{i+1}$ for any $i=1,\dots,l-1$.
 Every element of $\Gamma_{1,2}(S)$ may be expressed as follows:
@@
 (40) $p_{n}=\begin{cases}f_{n-1}(S)&\textup{$n$ is even},\\ f_{n}(S)&\textup{$n$ is odd},\end{cases}\quad q_{n}=\begin{cases}f_{n}(S)&\textup{$n$ is even},\\ f_{n-1}(S)&\textup{$n$ is odd}.\end{cases}$
-We may check it by the induction on $n$.
-By using this property, entries of $\Gamma_{1,2}(S)$ are closely related to the Chebychev polynomial.
-The following properties are well known for the Chebyshev polynomials.
+We may check it by the induction on $n$. By using this property, entries of $\Gamma_{1,2}(S)$ are closely related to the Chebyshev polynomials. The following properties are well known.
 Proposition 6.4 (e.g. [HM03, (1.4), (1.33b)]).
@@
 For the reader’s convenience, we give a proof following [FT19].
-When $20<r<$ 2, we may express $r=2\cos\theta$ for some $\theta\in\mathbb{R}$ with $0<\theta<\frac{\pi}{2}$.
+When $0<r<2$, we may express $r=2\cos\theta$ for some $\theta\in\mathbb{R}$ with $0<\theta<\frac{\pi}{2}$.
 By Proposition 6.3, $f_{n}(S)=qu_{n}(r)-pu_{n-1}(r)$ is a entry of an element in $\Gamma(S)$.
@@
 ∎
+For later, we show the following proposition.
 Proposition 6.6.
@@
 Let $p_{n}$ and $q_{n}$ be in (40).
-If $p_{n}$ or $q_{n}$ converges to some real number, then we have $qe^{\theta}-p=0$.
+If $p_{n}$ or $q_{n}$ converges to some real numbers, then we have $qe^{\theta}-p=0$.
 Moreover,
 (a) if $r=2$, then we have $p_{n}=q_{n}=p=q$ for any $n=0,1,2,\dots$.
-(b) if $r>2$, then both $p_{n}$ and $q_{n}$ converge to 00.
+(b) if $r>2$, then both $p_{n}$ and $q_{n}$ converge to 0.
 Proof.
@@
 By considering the maps $\gamma_{1}$ and $\gamma_{2}$, we may obtain $p_{n}=q_{n}=p=q$ for any $n\in\mathbb{Z}_{>0}$.
-If $r>2$, or equivalently, if $\theta>0$, it holds that $\sin n\theta>0$ for any $n\in\mathbb{Z}_{>0}$ and we have $\frac{\sinh(n+1)\theta}{\sinh n\theta}=\frac{e^{(n+1)\theta}-e^{-(n+1)\theta}}{e^{n\theta}-e^{-n\theta}}=\frac{e^{\theta}-e^{-(2n+1)\theta}}{1-e^{-2n\theta}}\to e^{\theta}$ when $n\to\infty$.
+If $r>2$, or equivalently, if $\theta>0$, it holds that $\sinh n\theta>0$ for any $n\in\mathbb{Z}_{>0}$ and we have $\frac{\sinh(n+1)\theta}{\sinh n\theta}=\frac{e^{(n+1)\theta}-e^{-(n+1)\theta}}{e^{n\theta}-e^{-n\theta}}=\frac{e^{\theta}-e^{-(2n+1)\theta}}{1-e^{-2n\theta}}\to e^{\theta}$ when $n\to\infty$.
 Since $u_{n}(r)=U_{n}(\cosh\theta)=\frac{\sinh(n+1)\theta}{\sinh n\theta}$, we have
@@
 $=\frac{\sinh n\theta}{\sinh\theta}\left(q\frac{\sinh(n+1)\theta}{\sinh n\theta}-p\right).$
-Unless $qe^{\theta}-p\neq 0$, it does not converge since $\sinh(2n+1)\theta\to\infty$ when $n\to\infty$. When $p=qe^{\theta}$, we have
+If $qe^{\theta}-p\neq 0$, it does not converge since $\sinh n\theta\to\infty$ when $n\to\infty$. When $p=qe^{\theta}$, we have
 (42) $f_{n}(S)$ $=q\frac{\sinh(n+1)\theta}{\sinh\theta}-qe^{\theta}\frac{\sinh n\theta}{\sinh\theta}$
@@
 It was shown in [FT19] by using the geometrical aspects of mutations.
-Here, we give a proof which is more similar to that in [BBH11], althoug we need some modifications for the real case.
+Here, we give an alternative proof which is more closer to that in [BBH11], althoug we need some modifications for the real case.
 In the previous section, for any $S\in\mathcal{S}^{+}$, $\Gamma(S)=\{\gamma_{t}(S)\}_{t}$ was defined as a collection indexed by $t$.
@@
   • (M2)
-    For presicely two indices $i$ in $\{1,2,3\}$, it holds that $S\leq\gamma_{i}(S)$.
+[…]
   • (M3)
@@
 (2) A triplet $S$ satisfies (M2) if and only if $2p^{2}>pqr\geq 2q^{2}$. In this case, $S$ satisfies $S\nleq\gamma_{1}(S)$ and $S\leq\gamma_{2}(S),\gamma_{3}(S)$.
-(3) A triplet $S$ satisfies (M3) if and only if $2q>pr$. In this case, $S$ satisfies $S\nleq\gamma_{1}(S),\gamma_{2}(S)$.
+[…]
 Proof.
@@
 Following a similar argument as in [BBH11, Lem. 2.1], we will prove some fundamental properties of these classes in $\mathcal{S}^{+}$.
+First, we show that the condition $(M3)$ does not appear in $\Gamma(S)$ for any cluster-positive $S$.
 Lemma 7.4 (see [BBH11, Lem. 2.1. (c)]).
@@
 Proof.
-Without loss of generality, we may assume $p\geq q\geq r>0$. By Proposition 7.3 (3), we have $2q>pr$ and $2p>qr$. By multiplying the corresponding sides, we have
+[…]
 (44) $4pq>pqr^{2}$
 and it implies $4>r^{2}$, namely, $r<2$. By Proposition 6.5, we may verify that $S$ is not cluster-positive. ∎
+By the above lemma, if $S$ is cluster-positive, then $\Gamma(S)$ consists of elements which satisfy either (M1) or (M2). In particular, (M1) gives some strong properties as follows.
 Lemma 7.5 (see [BBH11, Lem 2.1 (a)]).
 Let $S=(p,q,r)\in\mathcal{S}^{+}$ satisfy (M1). Then, the following statements hold.
-(a) We have three inequarities $p,q,r\geq 2$.
+(a) We have three inequalities $p,q,r\geq 2$.
 (b) Every element in $\Gamma(S)$ satisfies (M2) except for $S$. Here, we see $\Gamma(S)=\{\gamma_{t}(S)\mid t\}$ as a set.
@@
 (45) $S=S_{0}\overset{\gamma_{t_{1}}}{\mapsto}S_{1}\overset{\gamma_{t_{2}}}{\mapsto}S_{2}\overset{\gamma_{t_{3}}}{\mapsto}\cdots\overset{\gamma_{t_{k}}}{\mapsto}S_{k}=S^{\prime}$
-which satisfies $t_{i-1}\neq t_{i}$ ($i=1,2,\dots,k$), then we have $S_{i-1}\leq S_{i}$.
+which satisfies $t_{i-1}\neq t_{i}$ ($i=1,2,\dots,k$), we have $S_{i-1}\leq S_{i}$.
 In particular, for any $S^{\prime}\in\Gamma(S)$, there is a sequence (45) with $S_{i-1}\leq S_{i}$, and $S_{i-1}\neq S_{i}$ for any $i=1,2,\dots,k$.
-(d) This $S$ is the smallest element in $\Gamma(S)$ in terms of the partial order $\leq$. In particular, $S$ is cluster-positive.
+(d) This $S$ is the minimum element in $\Gamma(S)$ in terms of the partial order $\leq$. In particular, $S$ is cluster-positive.
 Proof.
@@
 (46) $pqr\geq 2p^{2},2q^{2},2r^{2}$
-For example, by multiplying both sides of two inequatries $pqr\geq 2p^{2}$ and $pqr\geq 2q^{2}$, we have $(pqr)^{2}\geq 4p^{2}q^{2}$, and it implies $r^{2}\geq 4$ and $r\geq 2$.
+For example, by multiplying both sides of two inequalities $pqr\geq 2p^{2}$ and $pqr\geq 2q^{2}$, we have $(pqr)^{2}\geq 4p^{2}q^{2}$, and it implies $r^{2}\geq 4$ and $r\geq 2$.
 Similarly, we have $p,q\geq 2$.
@@
 Let $S\in\mathcal{S}^{+}$ satisfy the condition (B) and let $\{S_{i}\}$ be a sequence of (47). Then, we have the following statements.
-(a) The sequence $\{S_{i}\}$ converges to $(2,2,2)$.
+(a) The sequence $\{S_{i}\}$ converges to (2,2,2)(2,2,2)( 2 , 2 , 2 ).
 (b) We have $C(S)=4$.
-The existence of the limit was shown in [BBH11, Thm. 5.1], but we can show that this limit is $(2,2,2)$.
+The existence of the limit was shown in [BBH11, Thm. 5.1], but we can show that this limit is (2,2,2)(2,2,2)( 2 , 2 , 2 ).
 Proof.
@@
 Now, we assume $a=b>2=c$.
-Then, we have $p_{i},q_{i}>2=r_{i}$ for any $i$, and we should choose the sequence of indices as $(1,2,1,2,\dots)$ or $(2,1,2,1,\dots)$.
-By Proposition 6.6, we have $p_{i}=q_{i}=a$ because $S_{i}$ converges to some nonzero number $a$.
+Then, we have $p_{i},q_{i}>r_{i}$ for enough large $i$, and we should choose the sequence of indices as (1,2,1,2,)(1,2,1,2$\dots$)( 1 , 2 , 1 , 2 , … ) or (2,1,2,1,)(2,1,2,1$\dots$)( 2 , 1 , 2 , 1 , … ).
+By Proposition 6.6, we have $p_{i}=q_{i}=a$ and $r_{i}=2$ because $S_{i}$ converges to some nonzero number $a$.
 However, $S_{i}=(a,a,2)$ satisfies (M1), and this contradicts the fact that $S_{i}$ satisfies (M2).
@@
 (b) Note that $C(S)=C(S_{i})=p_{i}^{2}+q_{i}^{2}+r_{i}^{2}-p_{i}q_{i}r_{i}$ does not depend on $i$.
-By considering the limit of $i\to\infty$, we have $C(S)=2^{2}+2^{2}+2^{2}-2\cdot 2\cdot 2=4=C(L(S))$.
+Moreover, the right hand side consists of the continuous function.
+Thus, by considering the limit $i\to\infty$, we have $C(S)=C(2,2,2)=4$.
 ∎
@@
 3.3].
-In particular, the proof is different when we consider $32<r<$ 3.
+In particular, the proof is different when we consider $2<r<3$.
 Proof.
-Since $S$ satisfies (M1), by Lemma 7.5 (a) and Proposition 7.3 (1), we have $\frac{1}{2}qr\geq p\geq q\geq r\geq 2$.
-Now, we fix $r\geq 2$ and we see $C(S)$ as a function of $(p,q)\in\mathbb{R}^{2}$ satisfying the above inequalities.
-The region where $(p,q)$ satisfies these inequalities is illustrated in Figure 3.
-Set $f(p,q)=p^{2}+q^{2}+r^{2}-pqr$.
-$p$$q$$qr=2p$$p=q$$r$$q$$(p_{1},q)$$(p,q)$$(p_{2},q)$
+[…]
+[…]
 Figure 3. The region of $(p,q)$
@@
 In this case, $p$ satisfies $p_{1}\leq p\leq p_{2}$ as in Figure 3.
-Then, since $f(p,q)$ is a quadratic function with the positive leading coefficient $1$, we have $f(p,q)\leq\max(f(p_{1},q),f(p_{2},q))$.
+Then, since $f(p,q)$ is a quadratic function with the positive leading coefficient 111, we have $f(p,q)\leq\max(f(p_{1},q),f(p_{2},q))$.
 We can show $f(p_{1},q)\geq f(p_{2},q)$ as follows:
@@
 $=(p_{1}-p_{2})(p_{1}-qr+p_{2})\geq 0.$
-The last inequality may be obtained by $p_{1}-p_{2}\leq 0$ and $p_{1}-qr+p_{2}=q-qr+\frac{1}{2}qr=-\frac{1}{2}q(r-2)\leq 0$.
-Thus, we have $f(p,q)\leq f(p_{1},q)=f(q,q)$.
-Moreover, since $f(q,q)=2q^{2}+r^{2}-q^{2}r=q^{2}(2-r)+r^{2}$ is a quadratic function of $q$ with the negative coefficient $02-r<$ 0 or a constant $r^{2}$, we have $f(q,q)\leq f(r,r)=C(r,r,r)$.
-(We used $q\geq r>0$.) Thus, we have $C(S)\leq C(r,r,r)$.
+[…]
 Next, we show Lemma 8.1. By Lemma 8.2 (1), it suffices to show that $C(r,r,r)=3r^{2}-r^{3}\leq 4$. It may be shown as follows:
@@
 Thus, $C(S)=4$ is equivalent to $p=q\geq r=2$. ∎
-9. The case (B)
-Recall that there are two classes (A) and (B) of cluster-positive elements. (See Definition 7.6.) By considering Lemma 7.7 and Lemma 8.2, we have the following theorem.
-Theorem 9.1.
-Let $S\in\mathcal{S}^{+}$ be cluster-positive.
-(a) This element $S$ satisfies (A) if and only if $C(S)<4$ or $S$ belongs to $\Gamma(p,p,2)$, $\Gamma(p,2,p)$, or $\Gamma(2,p,p)$ for some $p\geq 2$.
-(b) This element $S\in\mathcal{S}^{+}$ satisfies (B) if and only if $C(S)=4$ and $S$ does not belong to $\Gamma(p,p,2)$, $\Gamma(p,2,p)$, and $\Gamma(2,p,p)$ for any $p\geq 2$.
-Proof.
-By Lemma 7.7, if a cluster-positive element $S$ satisfies (B), then we have $C(S)=4$.
-Moreover, cluster-positive elements $S\in\mathcal{S}^{+}$ satisfying $C(S)=4$ and (A) should belong to $\Gamma(p,p,2)$, $\Gamma(p,2,p)$, or $\Gamma(2,p,p)$ by Lemma 8.2 (2).
-Thus, the claim holds.
-∎
-An element $S=(\frac{29}{2},\frac{19}{4},\frac{7}{2})$ would satisfy (B).
-In fact, I have already checked that there is a decreasing sequence $S_{0},S_{1},\dots,S_{50}$ by using a computer.
-However, I cannot show that this $S$ satisfies (B).
-(We may check $C(S)=4$, but it would be difficult to show $S\notin\Gamma(p,p,2)\cup\Gamma(p,2,p)\cup\Gamma(2,p,p)$ for any $p\geq 2$.)
-In [FT19, Rem. 4.7], another identification between (A) and (B) is shown.
 9. Integer skew-symmetrizable case
@@
 For this, we consider $\Gamma$ as a subgroup of $\mathrm{Aut}(\mathcal{M})$ generated by $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$.
-Recall that $\mathcal{M}$ is defined by the set of all $M=\left(\begin{smallmatrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{smallmatrix}\right)$ corresponding to a skew-symmetrizable matrix $B\in\mathrm{M}_{3}(\mathbb{R})$ as in Defnition 4.2.
+Recall that $\mathcal{M}$ is defined by the set of all $M=\left(\begin{smallmatrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{smallmatrix}\right)$ corresponding to a skew-symmetrizable matrix $B\in\mathrm{M}_{3}(\mathbb{R})$ as in Defnition 4.1.
 In this section, we consider the following subsets.
@@
 Assume $p,q,r\neq 0$.
-Since $p^{2},q^{2},r^{2}\in\mathbb{Z}$, we may express $p=l\sqrt{a}$, $q=m\sqrt{b}$, and $r=n\sqrt{c}$, where $\0l,m,n\in\mathbb{Z}\backslash\{0\}$ and $a,b,c\in\mathbb{Z}_{>0}$ do not have any square factors.
-Since $pqr=lmn\sqrt{abc}\in\mathbb{Z}$ and $\0lmn\in\mathbb{Z}\backslash\{0\}$, we need $\sqrt{abc}\in\mathbb{Z}$.
+Since $p^{2},q^{2},r^{2}\in\mathbb{Z}$, we may express $p=l\sqrt{a}$, $q=m\sqrt{b}$, and $r=n\sqrt{c}$, where $l,m,n\in\mathbb{Z}\backslash\{0\}$ and $a,b,c\in\mathbb{Z}_{>0}$ do not have any square factors.
+Since $pqr=lmn\sqrt{abc}\in\mathbb{Z}$ and $lmn\in\mathbb{Z}\backslash\{0\}$, we need $\sqrt{abc}\in\mathbb{Z}$.
 Let $p_{k}$ be the $k$th prime number, and consider factorizations $a=\prod p_{k}^{\varepsilon_{k}^{a}}$, $b=\prod p_{k}^{\varepsilon_{k}^{b}}$, and $c=\prod p_{k}^{\varepsilon_{k}^{c}}$.
-Since $a$, $b$, and $c$ do not have square factors, every exponent of $p_{k}$ should be 00 or $1$.
+Since $a$, $b$, and $c$ do not have square factors, every exponent of $p_{k}$ should be 0 or 111.
 Moreover, since $\sqrt{abc}=\sqrt{\prod p_{k}^{\varepsilon_{k}^{a}+\varepsilon_{k}^{b}+\varepsilon_{k}^{c}}}$ is an integer, we need $\varepsilon_{k}^{a}+\varepsilon_{k}^{b}+\varepsilon_{k}^{c}\in 2\mathbb{Z}$ for any $k$.
-In particular, $(\varepsilon_{k}^{a},\varepsilon_{k}^{b},\varepsilon_{k}^{c})$ is any of $(0,0,0)$, $(1,1,0)$, $(0,1,1)$, or $(1,0,1)$.
+In particular, $(\varepsilon_{k}^{a},\varepsilon_{k}^{b},\varepsilon_{k}^{c})$ is any of (0,0,0)(0,0,0)( 0 , 0 , 0 ), (1,1,0)(1,1,0)( 1 , 1 , 0 ), (0,1,1)(0,1,1)( 0 , 1 , 1 ), or (1,0,1)(1,0,1)( 1 , 0 , 1 ).
 Let
@@
 As Definition 7.6, every cluster-positive element $S\in\mathcal{S}$ satisfies the condition (A) or (B).
-However, when we consider $S\in\widehat{\mathcal{S}}$, the case (B) does not occur as follows.
+However, when we consider $S\in\widehat{\mathcal{S}}$, the case (B) does not occur.
 (This proposition is a generalization of [BBH11, Lem.
@@
 is a fundamental domain for the action of $\Gamma$ on $\widehat{\mathcal{S}}^{\mathrm{CP}}$ [BBH11, Thm. 5.1]. We generalize this result for the skew-symmetrizable case.
-For this, we introduce a partial order $\leq$ on $\mathcal{M}$ as each corresponding entry satisfies this order on $\mathbb{R}$.
-Then, this order is compatible for the order on $\mathcal{S}$ and the map $\mathrm{Sk}$.
+For this, we introduce a partial order $\leq$ on $\mathcal{M}$ as each corresponding entry satisfies this order on $\mathbb{R}$. Then, this order is compatible for the map $\mathrm{Sk}$.
 Lemma 9.4.
@@
 We show $M_{i-1}\leq M_{i}$.
-Consider a sequence defined by $S_{i}=\mathrm{Sk}(M_{i})$.
+Consider a sequence $S_{0}=\mathrm{Sk}(M_{0}),S_{1}=\mathrm{Sk}(M_{1}),\dots,S_{k}=\mathrm{Sk}(M_{k})$.
 Then, by Lemma 7.5 (c), we have $S_{i-1}\leq S_{i}$, and it implies that $M_{i-1}\leq M_{i}$ by Lemma 9.4.
 ∎
-Theorem 10.7.
+As a conclusion, we may give a fundamental domain for the action $\Gamma$. (The existence and the uniqueness of the minimum element was shown in [Sev13, Thm. 1.2].)
+Theorem 9.7 (cf. [Sev13, Thm. 1.2]).
 A set
 […]
-=$=\left\{\left(\begin{matrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{matrix}\right)\in{\mathcal{M}}_{\mathbb{Z}}^{+}\mid xyz\geq 2xx^{\prime},2yy^{\prime},2zz^{\prime}\right\}$
+=$=\left\{\left.\left(\begin{matrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{matrix}\right)\in{\mathcal{M}}_{\mathbb{Z}}^{+}\ \right|\ xyz\geq 2xx^{\prime},2yy^{\prime},2zz^{\prime}\right\}$
 is a fundamental domain for the action of $\Gamma$ on $\mathcal{M}_{\mathbb{Z}}^{\mathrm{CP}}$.
@@
 We show $\Gamma(M)\cap F\neq\emptyset$.
-By Proposition 10.3, there is $S=\gamma_{t}(\mathrm{Sk}(M))\in\Gamma(\mathrm{Sk}(M))$ satisfying (M1) for some finite sequence of indices.
+By Proposition 9.3, there is $S=\gamma_{t}(\mathrm{Sk}(M))\in\Gamma(\mathrm{Sk}(M))$ satisfying (M1) after some transformation $\gamma_{t}$.
 Then, since $\gamma_{t}(\mathrm{Sk}(M))=\mathrm{Sk}(\gamma_{t}(M))$, $\gamma_{t}(M)$ satisfies (M1).
@@
 Lemma 9.8.
-An element $M=\left(\begin{smallmatrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{smallmatrix}\right)\in\mathcal{M}^{+}$ is a fixed point for the action $\Gamma$ if and only if it holds that
+An element $M=\left(\begin{smallmatrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{smallmatrix}\right)\in\mathcal{M}^{+}$ is a fixed point for the action $\Gamma$ if and only if $\mathrm{Sk}(M)=(2,2,2)$, that is,
 (58) $xx^{\prime}=yy^{\prime}=zz^{\prime}=4.$
@@
 The equalities $\gamma_{1}(S^{\prime})=\gamma_{2}(S^{\prime})=S^{\prime}$ imply $p=q$ and $r=2$.
-Moreover, since $\gamma_{3}(S^{\prime})=(p,p,p^{2}-2)\leq S^{\prime}$, we need $p^{2}-2p\leq 2$.
+Moreover, since $\gamma_{3}(S^{\prime})=(p,p,p^{2}-2)\leq S^{\prime}$, we need $p^{2}-2\leq 2$.
 Since $p\geq 2$, the inequality $p^{2}-2\leq 2$ implies $p=2$.
 Thus, a maximum element $M^{\prime}$ should satisfy $\mathrm{Sk}(M^{\prime})=(2,2,2)$.
-By Lemma 10.8, $M^{\prime}$ must be expressed as in (60).
+By Lemma 9.8, $M^{\prime}$ is a fixed point and $M^{\prime}$ must be expressed as in (59).
 Moreover, since $\Gamma(M)=\Gamma(M^{\prime})=\{M^{\prime}\}$, $M=M^{\prime}$ holds.
@@
 9.3. Surjectivity of the Markov constants
-In [BBH11, Thm.
-1.2 (3)], they showed that there are no integral cluster-positive elements $S$ with $40<C(S)<$ 4.
-Moreover, an element $S$ satisfying (M1) and $C(S)=0$ is only $S=(3,3,3)$.
-On the other hand, by extending integer skew-symmetrizable matrices $\mathcal{M}_{\mathbb{Z}}^{\mathrm{CP}}$ (or $\widehat{\mathcal{S}}^{\mathrm{CP}}$), we can find some elements with $40\leq C(S)<$ 4 as in Example 10.18.
-More strongly, we obtain the following result.
+[…]
 Theorem 9.12.
@@
 Of course, this choice of $S$ is not unique. For example, we may take $S=(\sqrt{-n+9},\sqrt{-n+9},3)$ for $C(S)=n$.
-Next, we give all elements $S\in\widehat{\mathcal{S}}^{\mathrm{CP}}$ satisfying $40\leq C(S)<$ 4 and (M1). The classification in this region is important by the following property.
+Next, we give all elements $S\in\widehat{\mathcal{S}}^{\mathrm{CP}}$ satisfying $0\leq C(S)<4$ and (M1). The classification in this region is important by the following property.
 Proposition 9.14 ([BBH11, Lem. 3.3]).
@@
 Let $S=(p,q,r)\in{\mathcal{S}}^{+}$ satisfy (M1) and $p\geq q\geq r$ and let $C\in\mathbb{Z}_{<4}$.
-(a) There exists a real number $R>2$ satisfying $C(R,R,R)=C$.
+(a) There exists a real number $R>2$ satisfying $C(R,R,R)=C$. Moreover, this $R$ is unique.
 (b) If $C(S)\geq C$, then we have $r\leq R$, where $R$ is the real number satisfying $C(R,R,R)=C$.
@@
 Lemma 9.16.
-Let $S=(p,q,r)\in\mathcal{S}^{+}$ satisfy (M1) and $p\geq q\geq r$. Let $C\in\mathbb{Z}_{<4}$. If $C=C(S)$, then we have $2<r$ anf $p\leq r\sqrt{\frac{r^{2}-C}{r^{2}-4}}$.
+[…]
 Proof.
@@
 Example 9.18.
-Consider an element $S=(p,q,r)\in\widehat{\mathcal{S}}^{\mathrm{CP}}$ satisfying (M1) and $40\leq C(S)<$ 4.
+Consider an element $S=(p,q,r)\in\widehat{\mathcal{S}}^{\mathrm{CP}}$ satisfying (M1) and $0\leq C(S)<4$.
 The real number $R>2$ satisfying $C(R,R,R)=3R^{2}-R^{3}=0$ is $R=3$.
 Thus, by Lemma 9.15, under the assumption $p\geq q\geq r$, we need $2<r\leq 3$.
-Since $r^{2}\in\mathbb{Z}$, we have $\sqrt{5}\leq r\leq 3$.
+Since $r^{2}\in\mathbb{Z}$, we have $r=\sqrt{5},\sqrt{6},\sqrt{7},\sqrt{8},3$.
 Moreover, by setting $C=C(S)$, we have $p\leq r\sqrt{\frac{r^{2}-C}{r^{2}-4}}\leq 3\times\sqrt{\frac{3^{2}-0}{\sqrt{5}^{2}-4}}=9$ by Lemma 9.16.
 In particular, by considering $p^{2},q^{2},r^{2}\in\mathbb{Z}$, the number of possible $S$ are at most $5\times 81^{2}$.
-(By considering other conditions such as $pqr\in\mathbb{Z}$ and $r\leq q\leq p$, the number of elements which we need to check is much smaller than $5\times 81^{2}$.) We may write all cluster-positive orbits satisfying $40\leq C(S)<$ 4 as follows:
+(By considering other conditions such as $pqr\in\mathbb{Z}$ and $r\leq q\leq p$, the number of elements which we need to check is much smaller than $5\times 81^{2}$.) We may write all cluster-positive orbits satisfying $0\leq C(S)<4$ as follows:
 Proposition 9.19.
@@
 (e) $S$ is cluster-positive with $C(S)=4$ if and only if $S$ belongs to $\Gamma(p,p,2)$ for some $p\geq 2$ with $p^{2}\in\mathbb{Z}$ or their permutations.
+10. The case (B)
+Recall that there are two classes (A) and (B) among cluster-positive elements in $\mathcal{S}$.
+(See Definition 7.6.) For the skew-symmetric case, one identification was given by [FT19, Rem.
+4.6].
+Here, we slightly generalize it to the skew-symmetrizable case, and give another identification.
+Definition 10.1.
+For any cluster-positive element $M\in\mathcal{M}^{+}$, we define the class (A) and (B) by $\mathrm{Sk}(M)$ satisfies (A) and (B), respectively.
+Lemma 10.2.
+If a cluster-positive element $M\in M^{+}$ satisfies (B), the following statements hold.
+(a) There exists a unique infinite sequence
+(69) $M=M_{0}\overset{\gamma_{t_{1}}}{\mapsto}M_{1}\overset{\gamma_{t_{2}}}{\mapsto}M_{2}\overset{\gamma_{t_{3}}}{\mapsto}\cdots$
+such that $M_{i-1}\nleq M_{i}$ for any $i=1,2,\dots$.
+(b) This sequence has the limit $L(M)=\lim_{i\to\infty}M_{i}$.
+Moreover, $L(M)$ satisfies $\mathrm{Sk}(L(M))=(2,2,2)$.
+In particular, this limit $L(M)$ is a fixed point under the transformation $\Gamma$.
+Proof.
+Let $S=\mathrm{Sk}(M)$.
+Consider the sequence (47).
+Then, $S_{i-1}\nleq S_{i}$ holds.
+Set $M_{0}=M$ and $M_{i}=\gamma_{t_{i}}(M_{i-1})$.
+Then, by Lemma 9.4, this sequence also satisfies $M_{i-1}\nleq M_{i}$ as we desired.
+The uniqueness also follows from Lemma 9.4.
+(If there are another decreasing sequence, $S=\mathrm{Sk}(M)$ also has another decreasing sequence.) Thus, (a) holds.
+Next, we show (b).
+Let $M_{i}=\left(\begin{smallmatrix}a_{1;i}&a_{2;i}&a_{3;i}\\ a^{\prime}_{1;i}&a^{\prime}_{2;i}&a^{\prime}_{3;i}\end{smallmatrix}\right)$.
+By the assumption, each $a_{j;i}$ and $a^{\prime}_{j;i}$ ($j=1,2,3$) are monotonically decreasing for $i=0,1,2,\dots$.
+Moreover, since $M$ is cluster-positive, we have $a_{j;i},a^{\prime}_{j;i}\geq 0$.
+Thus, each component $a_{j;i},a^{\prime}_{j;i}$ should converge to some $u_{i},u^{\prime}_{i}\geq 0$, and we have $L(M)=\left(\begin{matrix}u_{1}&u_{2}&u_{3}\\ u^{\prime}_{1}&u^{\prime}_{2}&u^{\prime}_{3}\end{matrix}\right)$.
+Moreover, since $\mathrm{Sk}(S_{i})\overset{i\to\infty}{\longrightarrow}(2,2,2)$ by Lemma 7.7, we have $\sqrt{a_{j;i}a^{\prime}_{j;i}}\overset{i\to\infty}{\longrightarrow}2$.
+This implies $\sqrt{u_{j}u^{\prime}_{j}}=2$ ($j=1,2,3$) and $\mathrm{Sk}(L(M))=(2,2,2)$, which tells us that $L(M)$ is a fixed point by Lemma 9.8.
+∎
+By considering Lemma 7.7 and Lemma 8.2, we have the following theorem.
+Theorem 10.3.
+Let $M\in\mathcal{M}^{+}$ be cluster-positive element.
+(a) This element $M$ satisfies (A) if and only if $C(M)<4$ or $\mathrm{Sk}(M)$ belongs to $\Gamma(p,p,2)$, $\Gamma(p,2,p)$, or $\Gamma(2,p,p)$ for some $p\geq 2$.
+(b) This element $M$ satisfies (B) if and only if $C(M)=4$ and $\mathrm{Sk}(M)$ does not belong to $\Gamma(p,p,2)$, $\Gamma(p,2,p)$, and $\Gamma(2,p,p)$ for any $p\geq 2$.
+Moreover, in this case, there exists a decreasing sequence (69) uniquely, and it converges to some fixed point $L(M)$.
+Proof.
+Let $S=\mathrm{Sk}(M)$.
+By Lemma 7.7, if a cluster-positive element $S$ satisfies (B), then we have $C(S)=4$.
+Moreover, cluster-positive elements $S\in\mathcal{S}^{+}$ satisfying $C(S)=4$ and (A) should belong to $\Gamma(p,p,2)$, $\Gamma(p,2,p)$, or $\Gamma(2,p,p)$ by Lemma 8.2 (2).
+Thus, the claim holds.
+∎